### v2

# From Link Prediction to Forecasting: Addressing Challenges in Batch-based Temporal Graph Learning

Abstract
Dynamic link prediction is an important problem often considered in recent works proposing various approaches for learning temporal edge patterns. To assess their efficacy, models are evaluated on benchmark datasets involving continuous-time and discrete-time temporal graphs. However, as we show in this work, the suitability of common batch-oriented evaluation depends on the datasets’ characteristics, which can cause multiple issues: For continuous-time temporal graphs, fixed-size batches create time windows with different durations, resulting in an inconsistent dynamic link prediction task. For discrete-time temporal graphs, the sequence of batches can additionally introduce temporal dependencies that are not present in the data. In this work, we empirically show that this common evaluation approach leads to skewed model performance and hinders the fair comparison of methods. We mitigate this problem by reformulating dynamic link prediction as a link forecasting task that better accounts for temporal information present in the data.
Graph Neural Network, GNN, Temporal Graph, Dynamic Link Prediction, Dynamic Graph, Temporal Graph Learning, Dynamic Graph Learning, Temporal Graph Neural Network, TGNN, DyGNN, Dynamic Graph Neural Network
1 Introduction
[…]
Figure 1: Illustration of the issues in batch-based evaluation: (1) Information loss: continuous-time temporal edges are grouped into batches, causing some TGNNs to discard their temporal ordering. (2) Information leakage: discrete-time temporal edges that belong to the same snapshot are assigned to different batches, imposing an artificial ordering between edges that belong to the same snapshot. (3) Varying time window durations: using a fixed batch size creates time windows of different lengths when temporal edges are inhomogeneously distributed. (4) Tunable batch size: when different models choose different batch sizes, they group the temporal edges into different time windows, thus making their results incomparable because the specifics of their prediction tasks differ.
Many scientific fields study data that can be modelled as graphs, where nodes represent entities that are connected by edges. Examples include social (Lazer et al., 2009), financial (Bardoscia et al., 2021), as well as molecular networks (David et al., 2020). Apart from the mere topology of interactions, i.e., who is connected to whom, such data increasingly include information on when these interactions occurred. Depending on the temporal resolution, the resulting temporal graphs are often categorised as continuous-time or discrete-time (Kazemi et al., 2020): State-of-the-art data collection technology provides high-resolution continuous-time temporal graphs that capture each interaction’s exact (and possibly unique) occurrence time, leading to temporal graphs that, e.g., capture time-stamped online interactions (Kumar et al., 2019) or contacts between individuals via proximity (Vanhems et al., 2013). In contrast, discrete-time temporal graphs give rise to a temporally ordered sequence of static snapshots, where each snapshot contains interactions recorded within a (typically coarse-grained) time interval. Examples include political collaboration networks (Fowler, 2006) frequently available in yearly snapshots.
Building on the growing importance of temporal data and the success of graph neural networks (GNNs) for static graphs (Corso et al., 2024), deep graph learning has recently been extended to temporal (or dynamic) graphs (Feng et al., 2024). To this end, several temporal graph neural network (TGNN) architectures have been proposed that are able to simultaneously learn temporal and topological patterns. These architectures are often evaluated in dynamic link prediction, where the task is to predict the existence of edges during a future time window of length $\Delta t$, e.g., to provide recommendations to users (Kumar et al., 2019).
TGNNs utilise temporal batches to speed up training for dynamic link prediction. To construct these temporal batches, the sequence of temporally ordered edges is divided into a sequence of equally large chunks, i.e. containing the same number of edges, leading to several problems (cf. Figure 1): (1) Some TGNNs treat edges within a batch as if they occurred simultaneously, essentially discarding temporal information within a batch for continuous-time data. (2) In discrete-time temporal graphs, large snapshots containing edges with identical timestamps are often split into a sequence of batches. This introduces a temporal order not present in the data, thus leaking information that should not be available to the model because edges with the same timestamp end up in different batches. (3) Fixed-size batches in continuous-time temporal graphs are often associated with time windows of varying lengths ${\Delta t}$, effectively resulting in tasks of varying difficulty for each batch. (4) Related to the previous issue, treating the batch size as a tunable parameter affects the resulting window lengths and changes the task difficulty, e.g., from predicting at the minute to the hour level, thus making model performance between different models incomparable when they choose different batch sizes.
Addressing these important problems in the evaluation of temporal graph learning techniques, our work makes the following contributions:
  • We quantify the information loss and leakage due to the aggregation of edges into batches on 14 discrete-time and continuous-time temporal graphs showing how dynamic link prediction performance can be tuned by changing the batch size.
  • Replacing the current link prediction task, we formulate a link forecasting task based on time windows instead of batches to adequately consider the available temporal information.
  • We experimentally evaluate state-of-the-art TGNNs for link forecasting. Our results highlight substantial differences in model performance compared to the batch-based link prediction task. Our results suggest that memory-based TGNNs are ill-suited for discrete-time data, which has so far been overlooked due to the overestimation of model performance caused by information leakage in batch-based evaluation.
  • We provide implementations of the link forecasting task for commonly used graph learning frameworks such as DyGLib (Yu et al., 2023) and PyTorch Geometric (Fey & Lenssen, 2019).
Although batch-based processing is a technical necessity for efficient model training, we show that splitting edges into batches regardless of their occurrence time leads to many problems. With dynamic link forecasting, we propose a simple yet effective solution. While having similar runtime as dynamic link prediction, our time-window-based evaluation facilitates a fairer and more realistic evaluation that better reflects real-world scenarios.
2 Preliminaries and Related Work
Temporal Graphs.
A temporal, or dynamic graph $G=\left(V,E\right)$ is a tuple where $V$ is the set of $n=\left|V\right|$ nodes and $E$ is a chronologically ordered sequence of $m=|E|$ time-stamped edges defined as $E=\left\{(u_{1},v_{1},t_{1}),\dots,(u_{m},v_{m},t_{m})\right\}$ with $t_{1}\leq\dots\leq t_{m}$ (Wang et al., 2021b; Yu et al., 2023). Each node $v_{i}$ can have static node features $\mathbf{h_{i}}\in H_{V}$ and each edge $(u_{i},v_{j},t)$ can have edge features $\mathbf{e_{ij,t}}\in H_{E}$ that change over time. We assume that interactions occur instantaneously with timestamps $t\in\mathbb{N}$. Although timestamps $t\in\mathbb{N}$ are discrete, such temporal graphs are often categorized as continuous-time (Kazemi et al., 2020; Skarding et al., 2021). In contrast, discrete-time temporal graphs coarse-grain time-stamped edges into a sequence of static snapshots $\left(G_{t_{i}:t_{j}}\right)$, where $G_{t_{i}:t_{j}}=(V,E_{t_{i}:t_{j}})$ with $E_{t_{i}:t_{j}}=\left\{(u,v)\mid\exists(u,v,t)\in E:t_{i}\leq t\leq t_{j}\right\}$ (Xue et al., 2022).
Dynamic Link Prediction.
Given time-stamped edges up to time $t$, the goal of dynamic link prediction is to predict whether an edge $(v,u,t+1)$ exists at future time $t+1$ (Kazemi et al., 2020; Wang et al., 2021b; Poursafaei et al., 2022; Yu et al., 2023). In practice, it is often computationally infeasible to train and evaluate models on all possible edges one edge at a time. Thus, the chronologically ordered sequence of edges $E$ is usually divided into temporal batches $B^{+}_{i}$, where each batch has a fixed size of $b$ edges. Edges within the same batch are typically processed in parallel (Rossi et al., 2020; Su et al., 2024). In addition to the existing (positive) edges $(u,v,t)\in B^{+}_{i}$, non-existing (negative) edges $(u^{-},v^{-},t^{-})\in B_{i}^{-}$ are sampled and used for training and evaluation (see Appendix A for a brief overview of negative sampling approaches). This is done since real-world graphs are typically sparse and using all possible edges between all node pairs would lead to a large class imbalance and longer runtime. Formally, we define the dynamic link prediction task as follows:
Definition 2.1.
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $b$ be the batch size and $B_{i}^{+}:=\{(u_{j},v_{j},t_{j})\mid\exists(u_{j},v_{j},t_{j})\in E:b\cdot i\leq j<b\cdot(i+1)\}\}$ the set of $b$ edges in the $i$-th batch. We further use $B_{i}^{-}$ to denote a set of negative edges drawn using negative sampling as described in Appendix A. For a given batch $i$ we use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<t_{b\cdot i}\right\}$ to denote the past edges. The goal of dynamic link prediction is to find a model $f_{\theta}\bigl{(}u,v,t\mid\hat{E}_{i},H_{V},H_{\hat{E}_{i}}\bigr{)}$ with parameters $\theta$ that, for all edges $(u,v,t)\in B_{i}^{+}\cup B_{i}^{-}$ in each batch $i$, predicts whether $(u,v,t)\in B_{i}^{+}$ or $(u,v,t)\in B_{i}^{-}$.
State-of-the-Art TGNNs.
Current state-of-the-art dynamic link prediction methods, such as JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGN (Rossi et al., 2020) keep an up-to-date memory of temporal information in the graph by utilizing recurrent neural networks. TGAT extends graph attention to the temporal domain and replaces positional encodings in GAT (Velickovic et al., 2018) with a vector representation of time (Xu et al., 2020). TCL (Wang et al., 2021a) uses a transformer-based architecture to capture the nodes’ time-evolving properties. CAWN learns temporal motifs based on causal anonymous walks (Wang et al., 2021b). GraphMixer takes an attention-free and transformer-free approach, using an MLP-based link encoder, a mean-pooling-based node encoder, and an MLP-based link classifier for predictions (Cong et al., 2023). DyGFormer combines nodes’ historical co-occurrences as interaction targets of the same source node with a temporal patching approach to capture long-term histories (Yu et al., 2023). Further approaches for temporal graph learning exist, including DyGEM (Taheri et al., 2019), DySAT (Sankar et al., 2020), EvolveGCN (Pareja et al., 2020), and DBGNN (Qarkaxhija et al., 2022). For a recent survey of deep-learning-based dynamic link prediction, we refer to Feng et al. (2024).
Open Challenges in Dynamic Link Prediction
Training.
Recent works (Zhou et al., 2022, 2023; Feldman & Baskin, 2024; Su et al., 2024) identified issues in the training of memory-based TGNNs with large batch sizes: Processing edges belonging to the same batch in parallel ignores their temporal dependencies, resulting in varying performance depending on the chosen batch size. This issue, called temporal discontinuity, is caused by Problem (1) from above, i.e. treating edges within the same batch as if they occurred simultaneously (cf. Appendix D). To solve this issue, Zhou et al. (2023) introduce a distributed training framework using smaller batch sizes. The proposed solution PRES (Su et al., 2024) accounts for intra-batch temporal dependencies using a prediction-correction scheme. Feldman & Baskin (2024) propose a model with a decoupling strategy to mitigate this problem. However, these works focus on training instead of evaluation which is our work’s focus. Parallel to our work, Huang et al. (2024) present the unified temporal graph (UTG) representation, which facilitates the training of continuous-time graphs with discrete-time TGNNs. They also train continuous-time TGNNs on discrete-time graphs and, thereby, briefly mention information leakage as described in Problem (2). However, this work does not consider all Problems (1) – (4), nor does it include an empirical evaluation of how Problem (2) affects model performance, which is addressed in our work.
Evaluation.
Recent progress in terms of TGNN evaluation includes the temporal graph benchmark (TGB) (Huang et al., 2023) similar to the static open graph benchmark (OGB) (Hu et al., 2020). Poursafaei et al. (2022) identify problems with random negative sampling for dynamic link prediction and propose novel time-dependent negative sampling techniques to improve the evaluation of TGNNs. Gastinger et al. (2023) identify issues in the evaluation of temporal knowledge graph forecasting. Although none of the models used for this task overlap with regular TGNNs for dynamic link prediction, some of the problems can be related, e.g., differences in forecasting horizons leading to incomparable results.
Research Gap
The aforementioned issues in training and evaluating TGNNs for dynamic link prediction arise from applying standard procedures from static link prediction to temporal graph data. While recent works have identified how some of these issues affect the performance of some models in some data sets, we lack a systematic understanding of how the Problems (1) - (4) affect state-of-the-art temporal graph learning architectures across different types of continuous- and discrete-time temporal graph data. Moreover, we lack a better formulation of a dynamic link forecasting task, which mitigates these problems and can guide the development of new TGNNs.
3 From Link Prediction to Link Forecasting
Learning temporal patterns in a batch-oriented fashion leads to issues in continuous-time and discrete-time graphs. We demonstrate these issues in eight continuous-time and six discrete-time temporal graphs, whose characteristics are summarised in Table 1 and Appendix C. To mitigate these issues, we formulate the link forecasting task based on time windows with fixed duration.
Table 1: Characteristics of continuous (top) and discrete-time (bottom) temporal graphs. For each dataset, we list the number of nodes $n$ and edges $m$, the resolution of timestamps, the dataset’s total duration $T$, the average number of edges $\overline{|E_{t}|}$ with the same timestamp $t$, and the temporal density $\nicefrac{{T}}{{m}}$. The US Legis. dataset is measured in congresses (c). The other datasets are measured in conventional units of time,i.e. seconds (s), minutes (min), days (d) or years (a). See Appendix C for more information on each dataset.
Dataset $n$ $m$ Res. $T$ $\overline{|E_{t}|}$ $\nicefrac{{T}}{{m}}$
Enron 184 125 235 1 s 3.6 a 5.5 ± 16.6 908.2 s
UCI 1899 59 835 1 s 193.7 d 1.0 ± 0.3 279.7 s
MOOC. 7144 411 749 1 s 29.8 d 1.2 ± 0.5 6.2 s
Wiki. 9227 157 474 1 s 31.0 d 1.0 ± 0.2 17.0 s
LastFM 1980 1 293 103 1 s 4.3 a 1.0 ± 0.1 106.0 s
Myket 17 988 694 121 1 s 197.0 d 1.0 ± 0.0 24.5 s
Social 74 2 099 519 1 s 242.3 d 3.7 ± 2.5 10.0 s
Reddit 10 984 672 447 1 s 31.0 d 1.0 ± 0.1 4.0 s
UN V. 201 1 035 742 1 a 71.0 a 14 385.3 ± 7142.1 36.1 min
US L. 225 60 396 1 c 11.0 c 5033.0 ± 92.4 1.8 $\cdot 10^{-4}$ c
UN Tr. 255 507 497 1 a 31.0 a 15 859.3 ± 3830.8 32.1 min
Can. P. 734 74 478 1 a 13.0 a 5319.9 ± 1740.5 91.8 min
Flights 13 169 1 927 145 1 d 121.0 d 15 796.3 ± 4278.5 5.4 s
Cont. 692 2 426 279 5 min 28.0 d 300.9 ± 342.4 1.0 s
3.1 Problems in Batch-based Dynamic Link Prediction
Below, we show that batching causes information loss or leakage by ignoring the existing temporal order of links or inducing an artificial order between them, respectively. Temporal batches lead to inconsistent tasks because they produce prediction time windows of varying lengths when link densities vary over time. The batch-based evaluation can further lead to incomparable results between models when the batch size is treated as a hyperparameter and different models choose different batch sizes.
(1) Information Loss.
In Figure 2 we use normalised mutual information (NMI) (Vinh et al., 2010) to measure the information loss caused by splitting the temporal edges into batches. NMI quantifies how much information observing one random variable conveys about another random variable (see Appendix B for more information). It takes values between 00, meaning “no information”, and $1$, i.e. “full information”. By treating the index $i$ of each batch $B_{i}$ assigned to each edge $(u,v,t)\in B_{i}$ as one random variable and the associated edge’s timestamp $t$ as the other, we can measure the temporal information that is retained after dividing edges into batches. In this case, an NMI value of $1$ means that we can reconstruct the timestamps of edges correctly from their batch number, and a value of 00 means that batch numbers do not carry any information about timestamps, i.e. total loss of information.
In Figure 2(a), larger batches result in more information loss for continuous-time temporal graphs where timestamps have a high resolution. This is because assigning edges that occur at different times to the same batch discards their temporal ordering; the larger the batch size, the more information is lost. A batch size of $b=1$ preserves most temporal information and maximises NMI because we obtain a bijective mapping between almost all timestamps and batch numbers, except when multiple edges occur simultaneously.
Figure 2(b) shows the batch-size dependent NMI for discrete-time temporal graphs. The “optimal” batch size that retains most temporal information depends on the average number of links per snapshot and, thus, on the characteristics of the data. Too small batch sizes impose an ordering on the edges within the snapshots that is not present in the data while too large batches stretch across snapshots and discard the temporal ordering of edges from different batches.
(a) For continuous-time, the temporal ordering of edges within batches is discarded. With increasing batch size, more edges with different timestamps are assigned to the same time window, resulting in increased information loss.
(b) Small batches for discrete-time temporal graphs impose an ordering amongst the edges within snapshots that is not present in the data, which discards the information that edges of the same snapshot occur simultaneously. The NMI has its maximum near the average snapshot size (cf. Table 1) after which the values decrease again similar to continuous-time datasets.
Figure 2: NMI (y-axis) measures the temporal information loss as a function of different batch sizes (x-axis), where smaller NMI values indicate more information loss.
(2) Information Leakage.
The number of edges inside snapshots of discrete-time graphs is often larger than the batch size (cf. Section C.4). Snapshots are, therefore, divided into multiple batches containing edges with the same timestamp but are evaluated sequentially. The same problem can arise in continuous-time graphs in moments of high activity where the number of edges with a specific timestamp surpasses the batch size. Memory-based models utilise information from previous batches to make predictions for later batches by updating their internal node representations. Due to the issues above, this can include edges from the same snapshot. This leads to an unfair advantage for these models due to the information leakage of edge occurrence patterns in the already processed parts of the snapshot.
(3) Varying Time Window Durations.
In Figure 3, we show the temporal activity for a selection of temporal graphs, i.e. the number of time-stamped edges within a given time interval (see Section C.2 for the remaining datasets). The results show that many real-world temporal graphs have highly inhomogeneous activities. For batch-oriented evaluation, this introduces the issue that each fixed-size batch $B_{i}^{+}$ determines a time window with duration $t_{b\cdot(i+1)-1}-t_{b\cdot i}$, i.e., time windows are shorter during periods with higher activity and longer during periods with lower activity (more details in Section C.3).
Figure 3: Real-world datasets exhibit diverse edge occurrence patterns that are visualised using the edge density across time, i.e., histograms of three continuous- and one discrete-time temporal graphs counting the number of edges per timestamp. Dashed lines divide the datasets into 70% train, 15% validation, and 15% test sets as used in Section 4. The other datasets’ histograms are shown in Section C.2.
(4) Tunable Batch Size.
The varying amount of temporal information (see Figure 2) and the changing window durations (cf. Figure 3, 5, and 6) for different batch sizes highlight that changing the batch size influences the dynamic link prediction task. Effectively, the batch size is a hidden hyperparameter that directly impacts the task’s characteristics (and difficulty). This possibility of manipulating the task by changing the batch size makes the results incomparable between models with different batch sizes.
Consider the Myket dataset (Loghmani & Fazli, 2023) as an example: it contains users $v$ and Android applications $u$, connected at time $t$ when user $v$ installs application $u$. The timestamps are provided in seconds and edges occur roughly every 30s on average (cf. Table 1), making the expected time range for a batch with size $b=2$ approximately 30s. Thus, the task is to predict who installs which applications during this time window. Choosing $b=120$ or $b=2880$ turns the task into a prediction problem for approximately the next hour or day, respectively. For empirical validation, we compare the performance of state-of-the-art TGNNs using two different batch sizes during evaluation in Appendix D and show that the performance can vary substantially between different batch sizes.
3.2 Our Solution: Dynamic Link Forecasting
In real-world applications the prediction time window is inherently connected to the problem at hand, necessitating a task formulation that is chosen carefully for each dataset instead of for each model. To address the aforementioned issues, we propose link forecasting, reformulating link prediction as a task that utilizes a fixed prediction time window with a variable number of edges. Compared to dynamic link prediction with a fixed batch size, this task is both easier and harder: it is easier because it limits the window durations in which the temporal ordering of edges is lost and, additionally, does not introduce artificial temporal orderings that are not present in the data. It is harder because it prevents information leakage and ensures that the task cannot be tuned to fit the model.
Motivation.
The study of temporal information is at the centre of time series forecasting (Benidis et al., 2023) where the goal is to forecast the values of a time series for a given time window in the future. We define our task accordingly and interpret the temporal edges $E$ as $n^{2}$ Boolean time series, each of which takes the value 1 when an edge occurs. Standard multivariate models output a value for each timestamp over a forecasting horizon $h$. In large-scale temporal graphs, it is computationally infeasible to forecast the existence of all $n^{2}$ possible links, thus, only a sample of negative edges is considered instead. In continuous-time dynamic graphs, observations are available at high resolution, e.g. seconds, however, for many practical applications, predicting at lower granularity suffices. For example, it is typically enough to predict whether a customer purchases a certain product within the next day or week. Therefore, we consider forecasting for all timestamps $[t+1,t+h]$ during a time window at once instead of for each of them individually, and define the link forecasting task as follows:
Definition 3.1.
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $h$ be the time horizon and $W_{i}^{+}:=\{(u,v,t)\mid\exists(u,v,t)\in E:i\cdot h\leq t<(i+1)\cdot h\}$ the set of edges in the $i$-th time window. We use $W_{i}^{-}$ to denote a sample of $|W_{i}^{+}|$ negative edges that are sampled using one of the negative sampling approaches described in Appendix A and do not occur as positive edges in time window $[i\cdot h,(i+1)\cdot h)$. We further use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<i\cdot h\right\}$ to denote the set of past edges for time window $i$. The goal of dynamic link forecasting is to find a model $f_{\theta}(u,v,t|\hat{E}_{i},H_{V},H_{\hat{E}_{i}})$ with parameters $\theta$ that, for all edges $(u,v,t)\in W_{i}^{+}\cup W_{i}^{-}$ in each time window $i$, forecasts whether $(u,v,t)\in W_{i}^{+}$ or $(u,v,t)\in W_{i}^{-}$.
Crucially, this definition makes the evaluation independent of the batch size $b$ and instead introduces a time horizon $h$ that defines the forecasting time window and coarse-grains the temporal information. This requires deliberately choosing the temporal resolution based on the characteristics of each dataset and application (see Appendix E for examples). Similar to batch-based dynamic link prediction, adopting such a time-window-based perspective enables parallel processing of edges by aggregating temporal information. To summarize, link forecasting is based on a time horizon that affects and controls the amount of temporal information and ensures that the information is consistent for different models, facilitating fair performance comparisons.
Computational Cost
Let $m$ be the total number of links, where typically $m\ll n^{2}$ since real-world graphs are sparse, then we can assign all links to their corresponding time window in a single pass over the links in $O(m)$ by checking their interaction time. After each link is assigned, the time complexity during model evaluation is the same as for the batch-based approach. Since the number of links per time window varies and windows can become large during periods when many temporal edges occur, memory overflows cannot be precluded entirely. To mitigate this issue, time windows can be subdivided into smaller chunks for GPU-based computations. Since information leakage instead of information loss needs to be prevented between chunks of the same time window, steps such as memory updates need to be accumulated and done based on the whole time window. We report the runtime of both approaches for the experiments of Section 4 in Table 11.
Implementation
We provide implementations for our evaluation procedure in commonly used PyTorch libraries to simplify the adoption of our approach. Specifically, we implement a new DataLoader called SnapshotLoader [^1] https://github.com/M-Lampert/pytorch_geometric/blob/snapshot-loader/torch_geometric/loader/snapshot_loader.py that replaces the widely used TemporalDataLoader in PyTorch Geometric (Fey & Lenssen, 2019). We extend DyGLib (Yu et al., 2023) with a command line argument horizon that can be used in the evaluation pipeline [^2] https://github.com/M-Lampert/DyGLib. The latter was used for the experiments in this work and can be used to reproduce our results.
4 Link Prediction vs. Forecasting
We now experimentally evaluate the performance of nine state-of-the-art models (Kumar et al., 2019; Trivedi et al., 2019; Xu et al., 2020; Rossi et al., 2020; Wang et al., 2021a, b; Poursafaei et al., 2022; Cong et al., 2023; Yu et al., 2023), both for the (conventional) dynamic link prediction task as well as our proposed dynamic link forecasting task. We use implementation and model configurations provided by DyGLib (Yu et al., 2023) (cf. Appendix F) and repeat each experiment five times to obtain averages. We use historical negative sampling (Poursafaei et al., 2022) and train each model using batch-based training and validation with batch size $b=200$ which was found “to be a good trade-off between speed and update granularity” (Rossi et al., 2020) and adopted in later works (Poursafaei et al., 2022; Yu et al., 2023). Afterwards, we evaluate each trained model with weights obtained by batch-based training using both our proposed time-window-based evaluation method and the common batch-based evaluation approach.
Table 2: Average links per window $\overline{|W_{i}^{+}|}$ and standard deviation for horizon $h$ used in the evaluation. We chose $b=200$, i.e. $\overline{|B_{i}^{+}|}=200$, and appropriate horizons $h$ for each dataset to get $\overline{|W_{i}^{+}|}\approx 200$. NMI uses the time window and batch IDs of the test set (cf. Appendix B) to quantify how much the chosen chunks differ between the approaches.
Dataset $h$ $\overline{|W_{i}^{+}|}$ NMI
Enron 172 800s ($48$h) 214.1 ± 274.1 0.80
UCI 57 600s ($16$h) 208.5 ± 335.5 0.83
MOOC 1200s ($\nicefrac{{1}}{{3}}$h) 199.3 ± 167.2 0.88
Wiki. 3600s ($1$h) 211.7 ± 56.3 0.89
LastFM 21 600s ($6$h) 204.4 ± 120.2 0.91
Myket 5400s ($\nicefrac{{3}}{{2}}$h) 220.1 ± 133.6 0.91
Social 1800s ($\nicefrac{{1}}{{2}}$h) 186.1 ± 165.3 0.91
Reddit 900s ($\nicefrac{{1}}{{4}}$h) 226.0 ± 54.5 0.92
Continuous-Time Temporal Graphs.
Following previous work, we set the batch size to $b=200$ and choose the forecasting horizon $h$ such that the resulting time windows contain $200$ edges on average to facilitate comparability with the batch-oriented approach (see Table 2 for the exact values). We quantify the amount of information shared between the two approaches in Table 2 as the NMI score between time window ID and batch ID (cf. Appendix B for a detailed explanation).
Table 3: Test AUC-ROC scores for link forecasting and the relative performance change compared to link prediction for continuous-time graphs using the same trained models (standard deviations in Section G.2). The last row/column provides mean $\mu$ and standard deviation $\sigma$ of the absolute values (without sign) of the relative change per column/row.
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
Enron 84.0( $\uparrow$8.6%) 80.3( $\uparrow$9.2%) 67.9( $\downarrow$0.2%) 69.0( $\uparrow$17.5%) 75.7( $\uparrow$13.9%) 82.7( $\uparrow$3.6%) 75.1( $\uparrow$11.1%) 88.6( $\uparrow$9.0%) 84.5( $\uparrow$10.6%) 9.3%±5.1%
UCI 86.8( $\uparrow$4.2%) 60.2( $\uparrow$17.1%) 62.1( $\downarrow$1.5%) 55.2( $\downarrow$7.4%) 56.5( $\downarrow$3.0%) 72.5( $\uparrow$4.9%) 56.3( $\downarrow$6.2%) 80.2( $\downarrow$0.5%) 75.7( $\downarrow$0.6%) 5.0%±5.1%
MOOC 83.1( $\downarrow$2.1%) 79.0( $\downarrow$2.1%) 87.4( $\downarrow$1.2%) 79.9( $\downarrow$2.9%) 68.8( $\downarrow$2.2%) 59.8( $\downarrow$3.4%) 68.4( $\downarrow$5.8%) 70.3( $\downarrow$5.5%) 80.0( $\downarrow$1.5%) 3.0%±1.7%
Wiki. 81.5( $\downarrow$0.4%) 78.3( $\downarrow$0.1%) 83.7( $\downarrow$0.6%) 82.9( $\downarrow$0.7%) 71.3( $\downarrow$0.4%) 77.2( $\uparrow$0.1%) 84.6( $\downarrow$0.6%) 87.3( $\downarrow$0.6%) 79.8( $\downarrow$0.3%) 0.4%±0.2%
LastFM 76.3( $\downarrow$2.2%) 69.0( $\downarrow$3.7%) 79.2( $\downarrow$1.9%) 65.2( $\downarrow$4.7%) 66.3( $\downarrow$2.6%) 78.0( $\downarrow$0.2%) 62.5( $\downarrow$2.7%) 59.9( $\downarrow$9.2%) 78.2( $\downarrow$1.0%) 3.1%±2.6%
Myket 64.4( $\uparrow$0.6%) 64.1( $\downarrow$0.1%) 61.2( $\uparrow$0.1%) 57.8( $\uparrow$0.4%) 33.5( $\uparrow$3.1%) 52.6( $\uparrow$1.3%) 58.3( $\downarrow$0.3%) 59.8( $\uparrow$0.5%) 33.8( $\uparrow$3.0%) 1.0%±1.2%
Social 92.1( $\uparrow$0.8%) 92.2( $\downarrow$0.5%) 92.2( $\uparrow$0.5%) 92.5( $\downarrow$0.1%) 86.5( $\downarrow$1.4%) 84.9( $\downarrow$1.1%) 94.7( $\downarrow$0.6%) 94.6( $\uparrow$0.6%) 97.3( $\uparrow$0.0%) 0.6%±0.4%
Reddit 80.6( $\downarrow$0.0%) 79.5( $\uparrow$0.0%) 80.4( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 80.2( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 76.2( $\downarrow$0.1%) 77.1( $\downarrow$0.1%) 80.2( $\uparrow$0.0%) 0.0%±0.1%
$\mu\pm\sigma$ 2.4%±2.9% 4.1%±6.1% 0.8%±0.7% 4.2%±6.0% 3.3%±4.4% 1.8%±1.9% 3.4%±4.0% 3.2%±4.0% 2.1%±3.6%
Table 3 shows AUC-ROC scores for time-window-based link forecasting and the relative change compared to the batch-based evaluation of dynamic link prediction (average precision scores are provided in Section G.3). The change in performance between our window-based and the batch-based approach largely depends on the dataset: Datasets with a similar window duration for all fixed-sized batches (quantified by NMI scores close to one in Table 2), such as Wikipedia, Reddit, or Myket, only exhibit small performance differences. This is expected since we chose the horizon $h$ to produce time windows close to the same average size as the fixed-sized batches. Nevertheless, we observe lower NMI values in Table 2 for datasets with inhomogeneously distributed temporal activity such as Enron or UCI – i.e. the time windows do not fit the fixed-sized batches well. These datasets with lower NMIs show substantial performance changes across models. This highlights that the performance scores of batch-based evaluation are skewed and may not reflect the models’ performance in a real-world setting on temporal datasets with inhomogeneous activities.
Figure 4: Model performance of selected continuous-time datasets varies substantially over time which is visualised by the AUC-ROC scores for each batch and time window individually. For both datasets, the best-performing memory-based and non-memory-based model is shown. The time (x-axis) is limited to the test set. The scores for other models and datasets are reported in Section G.4.
To investigate the causes of the performance differences of both evaluation approaches, we analyse the AUC-ROC scores for each batch and time window individually in Figure 4 (and Section G.4 for other datasets and models). We observe that the performance between both approaches diverges most during phases of irregular temporal activity. Specifically, the batch-based evaluation is affected by an anomaly in the Enron dataset at day 1181, where the performance of all models drops substantially. At this time, 1705 temporal edges with identical timestamps are evaluated in 9 consecutive batches. In contrast, our approach uses a single time window for all 1705 edges with identical time stamps. During the second half of the observation period of the Enron dataset, our approach leads to more time windows compared to the number of batches. For our window-based approach, most models perform better during this period compared to their performance in the first half. This increase in performance during the second half of the observation period is masked by the batch-based evaluation, which aggregates this long period of time in a very small number of batches. A similar (daily) pattern can be observed for the MOOC dataset, where each period of high activity is followed by a period of low activity. Our time-window-based evaluation reveals that most models are particularly unreliable during the low activity periods which is again masked by the batch-based approach.
Table 4: Test AUC-ROC scores as in Table 3 for discrete-time graphs.
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
UN V. 54.0( $\downarrow$26.7%) 52.2( $\downarrow$28.2%) 51.3( $\downarrow$27.1%) 54.4( $\uparrow$3.0%) 53.7( $\uparrow$7.1%) 89.6( $\uparrow$0.0%) 53.4( $\uparrow$0.6%) 56.9( $\uparrow$1.1%) 65.2( $\uparrow$3.5%) 10.8%±12.6%
US L. 52.5( $\downarrow$6.8%) 61.8( $\downarrow$22.6%) 57.7( $\downarrow$31.2%) 78.6( $\uparrow$0.2%) 82.0( $\uparrow$0.2%) 68.4( $\uparrow$1.3%) 75.4( $\downarrow$0.3%) 90.4( $\uparrow$0.2%) 89.4( $\uparrow$0.0%) 7.0%±11.7%
UN Tr. 57.7( $\downarrow$12.8%) 50.3( $\downarrow$20.4%) 54.3( $\downarrow$14.0%) 64.0( $\uparrow$3.6%) 67.6( $\uparrow$4.5%) 85.6( $\downarrow$1.0%) 63.7( $\uparrow$4.5%) 68.6( $\uparrow$3.4%) 70.7( $\uparrow$3.4%) 7.5%±6.6%
Can. P. 63.6( $\downarrow$0.5%) 67.5( $\uparrow$1.2%) 73.2( $\downarrow$0.2%) 72.7( $\uparrow$1.5%) 70.0( $\uparrow$2.9%) 63.2( $\uparrow$0.4%) 69.5( $\uparrow$2.0%) 80.7( $\downarrow$0.6%) 85.5( $\downarrow$12.5%) 2.4%±3.9%
Flights 67.4( $\downarrow$3.1%) 66.0( $\downarrow$4.3%) 68.1( $\downarrow$1.0%) 72.6( $\uparrow$0.0%) 66.0( $\uparrow$0.1%) 74.6( $\uparrow$0.0%) 70.6( $\downarrow$0.0%) 70.7( $\downarrow$0.0%) 68.8( $\uparrow$0.1%) 1.0%±1.6%
Cont. 95.6( $\uparrow$0.1%) 94.9( $\downarrow$0.6%) 96.6( $\uparrow$0.5%) 95.9( $\uparrow$0.6%) 86.7( $\uparrow$4.1%) 93.0( $\uparrow$0.9%) 95.7( $\uparrow$1.3%) 95.2( $\uparrow$1.1%) 97.7( $\uparrow$0.6%) 1.1%±1.2%
$\mu\pm\sigma$ 8.3%±10.2% 12.9%±12.2% 12.3%±14.1% 1.5%±1.5% 3.1%±2.7% 0.6%±0.5% 1.5%±1.7% 1.1%±1.2% 3.4%±4.8%
Discrete-Time Temporal Graphs.
We set $h=1$ to obtain one time window per snapshot, predicting for time intervals ranging from five minutes to a year, depending on the datasets (cf. Table 1). Similar to Table 3, Table 4 shows the AUC-ROC scores and highlights performance differences between both evaluation approaches for discrete-time temporal graphs. For link forecasting, the performance of memory-based models (JODIE, DyRep, TGN) tends to decrease more than for other methods. This is expected since these models incorporate information about the present snapshot by updating their memory based on prior batches, which means using part of the snapshot’s edges to predict its remaining edges — this happens when the batch size is smaller than the snapshot, thus splitting the snapshot into a sequence of batches that are processed sequentially. We designed our evaluation method to prevent this information leakage, which explains the substantial drop in performance. We provide further results showing that our approach prevents information leakage in Section G.5.
Realistic Evaluation.
In the evaluation above, we orient the time horizon $h$ on the batch size $b$ to achieve as much comparability as possible between the link forecasting and prediction tasks. Thereby, we emphasize performance differences between both approaches since we use the same trained models and “only” group the edges differently. In real-world applications, however, the models’ performance should be evaluated using time horizons guided by the specific use case or research question. We provide examples for such an evaluation using the above datasets in Appendix E.
5 Conclusion
In this work, we considered issues in current evaluation practices for dynamic link prediction in temporal graphs. To address computational limitations, edges are split into fixed-size batches, making the task “too easy” but at the same time “too hard”: “Too easy” because state-of-the-art approaches for dynamic link prediction have treated the batch size as a tunable parameter. Changing the batch size, however, changes the task difficulty, resulting in incomparable results between models using different batch sizes. In discrete-time temporal graphs, batches that include edges from the same timestamp cause information leakage that can be utilised by memory-based TGNNs. “Too hard” since in continuous-time temporal graphs, fixed-size batches create varying-length time windows leading to tasks of different difficulty for each duration, requiring the model to capture multiple temporal resolutions at once. Furthermore, for edges within a batch, information regarding their temporal ordering is lost. In discrete-time temporal graphs where snapshots are typically larger than the batch size, batches additionally impose an ordering of edges not present in the data.
We solve these issues by formulating the dynamic link forecasting task. Dynamic link forecasting acknowledges the resolution at which temporal interaction data is recorded and explicitly considers a forecasting horizon corresponding to a time window of fixed duration. Depending on the dataset and problem setting, the horizon may span minutes, hours, or longer, but crucially, time windows always span the same length. We evaluated the dynamic link forecasting performance of nine state-of-the-art temporal graph learning approaches on 14 real-world datasets, comparing it to dynamic link prediction. We found substantial differences, especially for memory-based TGNNs. We provide implementations for PyTorch Geometric and DyGLib to facilitate practical applications of our evaluation approach.
Limitations and Open Issues
Limitations of our work include that our reformulation of the dynamic link prediction task suggests time-window-based approaches for model training, which however goes beyond the scope of our paper. Furthermore, the metrics used in this work consider the problem of dynamic link prediction or forecasting as binary classification, using one negative sample for each positive edge. Other approaches (You et al., 2022) consider the problem as a ranking task (e.g. using the mean reciprocal rank (MRR)) and compare each positive edge against a large number of negative samples. These approaches typically use only one positive sample per batch which alleviates the information loss and provides a better estimate of the models’ precision due to the large number of negative samples. In contrast, our window-based approach allows the parallel processing of many positive samples in each time window, leading to a considerably faster evaluation by reasonably simplifying the task. Specifically, our simplifying assumption is that it is enough to make predictions at a lower resolution than available in the data, resulting in a trade-off between evaluation runtime and granularity of temporal information available to the TGNN models. For example, it may not be required to predict whether a customer will purchase a certain product within the next second; making such a prediction for the next day or week may be sufficient.
Impact Statement
There are no expected negative societal impacts that go beyond those of other foundational works in machine learning research.
Acknowledgements
Moritz Lampert and Christopher Blöcker acknowledge funding from the German Federal Ministry of Education and Research, Grant No. 100582863 (TissueNet). Ingo Scholtes and Christopher Blöcker acknowledge funding through the Swiss National Science Foundation, Grant No. 176938.
Appendix A Negative Sampling Approaches
Dynamic link prediction is typically framed as a binary classification problem to predict class 1 for existing links during a certain time window and 0 otherwise. Due to the sparsity of most real-world graphs, it usually suffices to train and evaluate using all existing (positive) edges and a sample of non-existing (negative) edges out of all possible edges $V^{2}$. In static link prediction, negative edges are typically sampled randomly from $V^{2}$ without replacement but Poursafaei et al. (2022) showed that this technique is ill-suited for dynamic link prediction. One reason is rooted in the characteristics of temporal graphs where already-seen interactions tend to repeat several times during the observation period. To address this issue, Poursafaei et al. (2022) introduced negative sampling which we cover in the following.
Given a training set $E_{\text{train}}$ and test set $E_{\text{test}}$, each containing a sequence of edges $E_{t_{\text{min}}:t_{\text{max}}}$ in the temporal graph $G$, we can define the following commonly used sampling strategies for drawing negative samples $B_{i}^{-}$ for batch $B^{+}_{i}$ with $|B^{+}_{i}|=|B_{i}^{-}|$ (Poursafaei et al., 2022; Yu et al., 2023).
  • Random: Sample $B_{i}^{-}$ from $V^{2}$ without replacement. The subgraph corresponding to $B^{+}_{i}$ is assumed to be sparse, making it unlikely to sample a positive edge $e\in B_{i}$ as negative.
  • Historic: Sample $B_{i}^{-}$ without replacement from all training edges $E_{\text{hist}}=E_{\text{train}}\setminus\left\{(v_{j},u_{j})|t_{b\cdot i}\leq t_{j}\leq t_{b\cdot(i+1)}\right\}$ except the ones appearing at the same time as the edges in $B^{+}_{i}$. If $|E_{\text{hist}}|<|B^{+}_{i}|$, draw the remaining edges randomly as described above.
  • Inductive: Sample $B_{i}^{-}$ without replacement from all unseen test edges $E_{\text{ind}}=E_{\text{test}}\setminus(E_{\text{train}}\cup\left\{(v_{j},u_{j})|t_{b\cdot i}\leq t_{j}\leq t_{(i+1)\cdot b}\right\})$ except the ones appearing at the same time as the edges in $B^{+}_{i}$. If $|E_{\text{ind}}|<|B^{+}_{i}|$, draw the remaining edges randomly as described above.
Note that we leave out the validation set $E_{\text{val}}$ for simplicity. Negative edges for $E_{\text{val}}$ can be sampled as for $E_{\text{test}}$.
Appendix B Normalized Mutual Information
Normalized mutual information is an information-theoretic measure based on mutual information. It is based on mutual information, which for two random variables $X$ and $Y$ captures the bits of information we gain about the outcome of $Y$ if we know the outcome of $X$ and vice-versa. A formal definition is given in the following:
Definition B.1 (Mutual Information).
Consider two random variables, $X$ and $Y$ with joint probability mass function $p\left(x,y\right)$ and marginal probability mass functions $p\left(x\right)$ and $p\left(y\right)$. Mutual information is the reduction in the uncertainty of $X$ due to the knowledge of $Y$ defined as (Cover & Thomas, 2006)
$I\left(X,Y\right)=\sum_{x\in X}\sum_{y\in Y}p\left(x,y\right)\log\frac{p\left(x,y\right)}{p\left(x\right)p\left(y\right)}.$
Note that mutual information can be defined using different logarithms. The intuitive understanding described above using bits of information is defined using $\log_{2}$. This work utilizes the implementation of the Python library scikit-learn (Pedregosa et al., 2011) which uses the natural logarithm $\log_{e}$.
The specific value of mutual information depends on the entropy
$H(X)=-\sum_{x\in X}p(x)\log p(x)$
of the underlying random variables, and is thus difficult to compare across different settings. To address this issue, normalized mutual information provides a measure between zero and one that is normalized based on the entropies of the underlying random variables (Vinh et al., 2010). We use the following normalization as implemented in scikit-learn’s function normalized_mutual_info_score:
$\text{NMI}(X,Y)=\frac{I(X,Y)}{\frac{1}{2}\cdot(H(X)+H(Y))}$
In the context of our work, we use the NMI to capture the loss of temporal information. In Figure 2, we use the NMI to measure how much information about the edges’ timestamps is lost by grouping the edges into batches. Specifically, we use the number $i$ of each batch $B_{i}^{+}\cup B_{i}^{-}$ of the definition of dynamic link prediction in Section 2 assigned to each edge in the corresponding batch as one random variable and the timestamps of the edges as the other. With this setup, we can measure how much information about the timestamps of edges we gain – or keep – based on the batch number only.
Additionally, we use the NMI in Table 2 to quantify the difference between the assignments of edges to time windows and batches respectively. Similar as above, we assign to each edge in the test set of each dataset its batch number and also a time window number corresponding to the time window the edge belongs to and then compute the NMI between those two variables.
In the following, we present an educational example that illustrates the usage of NMI in the context of our work to facilitate comprehension among readers who may not be familiar with the measure:
Example B.1.
Let $G$ be a temporal graph with nodes $V=\left\{a,b,c\right\}$ and temporal edges
$E=\left((a,b,1),(b,c,2),(c,a,2),(a,c,4),(a,b,5),(b,a,5)\right).$
We define the timestamps $t_{i}$ of the edges $e_{i}\in E$, the batch number for $b=2$ and the window number with $h=1$ as random variables $X,Y$ and $Z$, respectively. We obtain the following realizations of $X,Y$ and $Z$:
$\mathbf{x}$ $=(1,2,2,4,5,5)$
$\mathbf{y}$ $=(1,1,2,2,3,3)$
$\mathbf{z}$ $=(1,2,2,3,4,4)$
In the next step, we compute the marginal probability mass functions $p(X=x),p(Y=y)$ and $p(Z=z)$:
$p(X=1)$ $=\frac{1}{6},$ $p(X=2)$ $=\frac{2}{6},$ $p(X=4)$ $=\frac{1}{6},$ $p(X=5)$ $=\frac{2}{6},$
$p(Y=1)$ $=\frac{2}{6},$ $p(Y=2)$ $=\frac{2}{6},$ $p(Y=3)$ $=\frac{2}{6},$
$p(Z=1)$ $=\frac{1}{6},$ $p(Z=2)$ $=\frac{2}{6},$ $p(Z=3)$ $=\frac{1}{6},$ $p(Z=4)$ $=\frac{2}{6}.$
Afterward, we compute the joint probability mass functions $p(X=x,Y=y)$, $p(X=x,Z=z)$, and $p(Y=y,Z=z)$:
$p(X=1,Y=1)$ $=\frac{1}{6},$ $p(X=2,Y=1)$ $=\frac{1}{6},$$\dots$
$p(X=1,Z=1)$ $=\frac{1}{6},$ $p(X=2,Z=2)$ $=\frac{2}{6},$$\dots$
$p(Y=1,Z=1)$ $=\frac{1}{6},$ $p(Y=1,Z=2)$ $=\frac{1}{6},$$\dots$
Using these marginal and joint probabilities, we compute the mutual information $I(X,Y)$, $I(X,Z)$, and $I(Y,Z)$ using the formula from above:
$I(X,Y)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{1}{6}\cdot\frac{2}{6}}+\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 0.868$
$I(X,Z)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{1}{6}\cdot\frac{1}{6}}+\frac{2}{6}\log_{e}\frac{\frac{2}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 1.330$
$I(Y,Z)$ $=\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{1}{6}}+\frac{1}{6}\log_{e}\frac{\frac{1}{6}}{\frac{2}{6}\cdot\frac{2}{6}}+\dots\approx 0.868$
Finally, we obtain the normalized mutual information (NMI) by normalizing based on the entropies of both random variables:
$\text{NMI}(X,Y)$ $=\frac{0.868}{\frac{1}{2}(1.330+1.099)}\approx 0.715$
$\text{NMI}(X,Z)$ $=\frac{1.330}{\frac{1}{2}(1.330+1.330)}\approx 1.0$
$\text{NMI}(Y,Z)$ $=\frac{0.868}{\frac{1}{2}(1.099+1.330)}\approx 0.715$
Note that the NMI calculations in the main part of the paper are done accordingly. Specifically, the NMI values of the bottom-most continuous-time dataset and the right-most discrete-time dataset in Figure 1 are calculated as demonstrated with $\text{NMI}(X,Z)$, while the other NMI values in the figure as well as all values in Figure 2 are calculated following the example of $\text{NMI}(X,Y)$. Table 2 reports NMI values that follow the example calculation of $\text{NMI}(Y,Z)$.
Appendix C Datasets
In this work, we use eight continuous-time and six discrete-time datasets, listed in Table 1. In the following, we describe what systems were observed to create the datasets and analyse the data in more detail.
C.1 Descriptions
We list details about each dataset in the following:
  • Enron (Shetty & Adibi, 2004) is a bipartite continuous-time graph where nodes are users and the temporal edges represent emails sent between users. Emails with multiple recipients are recorded as separate and simultaneously occurring edges, one per recipient. The temporal edges are resolved at the second level and the dataset spans approximately 3.6 years.
  • UCI (Panzarasa et al., 2009) is a unipartite continuous-time social network dataset from an online platform at the University of California at Irvine. The nodes represent students and the timestamped edges represent communication between the students. The dataset spans approximately six and a half months.
  • MOOC (massive open online course) (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent users and units in an online course, such as problems or videos. Temporal edges are resolved at the second level and encode when a user interacts with a unit of the online course. The dataset spans approximately one month.
  • Wikipedia (Wiki.) (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent editors and Wikipedia articles. The timestamped edges are resolved at the second level and represent when an editor has edited an article. The dataset spans approximately one month.
  • LastFM is a bipartite continuous-time graph where nodes represent users and songs. Temporal edges are resolved at the second level and model the users’ listening behavior and represent when a user has listened to a song. The dataset was originally published by Celma (2010) and later filtered by Kumar et al. (2019) for use in a temporal graph learning context.
  • Myket (Loghmani & Fazli, 2023) is a bipartite continuous-time graph where nodes represent users and Android applications. The timestamped edges represent when a user installed an application. The dataset spans approximately six and a half months.
  • Social Evolution (Social) (Madan et al., 2012) is a unipartite continuous-time graph of the proximity between the students in a dormitory, collected between October 2008 and May 2009 using mobile phones. Temporal edges connect students when they are in proximity and are resolved in seconds.
  • Reddit (Kumar et al., 2019) is a bipartite continuous-time graph where nodes represent Reddit users and their posts. The timestamped edges are resolved in seconds and represent when a user has made a post on Reddit. The dataset spans approximately one month.
  • UN Vote (UN V.)(Voeten et al., 2009; Poursafaei et al., 2022) is a weighted unipartite discrete-time graph of votes in the United Nations General Assembly between 1946 and 2020. Nodes represent countries and edges connect countries if they both vote “yes”. The dataset is resolved at the year level and edge weights represent how many times the two connected countries have both voted “yes” in the same vote.
  • US Legislators (US L.) (Huang et al., 2020; Fowler, 2006; Poursafaei et al., 2022) is a weighted unipartite discrete-time graph of interactions between legislators in the US Senate. Nodes represent legislators and edges represent co-sponsorship, i.e., edges connect legislators who co-sponsor the same bill. The dataset is resolved at the congress level and edge weights encode the number of co-sponsorships during a congress.
  • UN Trade (UN Tr.) (MacDonald et al., 2015; Poursafaei et al., 2022) is a directed and weighted unipartite discrete-time graph of food and agricultural trade between countries where nodes represent countries. The dataset spans 30 years and is resolved at the year level. Weighted edges encode the sum of normalized agriculture imports or exports between two countries during a given year.
  • Canadian Parliament (Can. P.) (Huang et al., 2020; Poursafaei et al., 2022) is a weighted unipartite discrete-time political network where nodes represent Members of the Canadian Parliament (MPs) and an edge between two MPs means that they have both voted “yes” on a bill. The dataset is resolved at the year level and the edges’ weights represent how often the two connected MPs voted “yes” on the same bill during a year.
  • Flights (Schäfer et al., 2014; Poursafaei et al., 2022) is a directed and weighted unipartite discrete-time graph where nodes represent airports and edges represent flights during the COVID-19 pandemic. The edges are resolved at the day level and their weights are given by the number of flights between two airports during the respective day.
  • Contacts (Cont.) (Sapiezynski et al., 2019; Poursafaei et al., 2022) is a unipartite discrete-time proximity network between university students. Nodes represent students who are connected by an edge if they were in close proximity during a time window. The dataset is resolved at the 5-minute level and spans one month.
C.2 Edge Activity patterns
Similar to Figure 3, we show the link densities over time for all datasets in Figure 5. While some datasets such as Enron, UCI, and LastFM contain very inhomogeneously distributed activity, others exhibit reoccurring or seasonal patterns, e.g. MOOC, Wikipedia, and Reddit. In addition to seasonality, we also observe trends like a growing number of edges in the Myket, UN Vote, and UN Trade datasets, which are also often found in standard time series data (Brockwell & Davis, 2016).
(a) Continuous-time temporal graphs resolved in seconds. For visualisation purposes, edges are binned into 6-hour time periods and then represented as one bar in the histogram.
(b) Discrete-time temporal graphs where each bar in the histogram corresponds to one snapshot. Since the snapshots capture different time ranges for each dataset (see Table 1), each dataset has its own temporal resolution.
Figure 5: Real-world datasets exhibit diverse edge occurrence patterns that are visualised using the edge density across time, i.e., histograms counting the number of edges per timestamp. Dashed lines divide the datasets into 70% train, 15% validation, and 15% test sets as used in Section 4.
C.3 Time Window Durations for Different Batch Sizes
In Figure 6 we evaluate the dependency between batch size and time window for empirical temporal graphs. The figure shows the distribution of batch durations (y-axis) for different batch sizes (x-axis). We observe that, both in continuous- and discrete-time temporal graphs, a single batch size can create time windows with varying durations even within the same dataset. For continuous-time temporal graphs, we typically have much bigger batches than edges per timestamp such that the time window defined by the batches becomes long (cf. Table 1). The number of edges per snapshot in discrete-time temporal graphs is generally larger than the batch size $b$ in any period regardless of the density (Table 1). This means that edges in a batch often belong to the same snapshot leading to small window durations.
(a) Continuous-time temporal graphs: Batch size $b$ determines the average time window length. However, a single batch size creates time windows with various lengths within and across datasets.
(b) Discrete-time temporal graphs: Fixed-size batches fall mostly within snapshots when the batches are much smaller than the snapshots. Depending on the dataset, larger batches can also span across many snapshots.
Figure 6: Using a low opacity value for individual points, the distribution of time window durations is shown for different batch sizes. I.e. points appear less see-through with an increasing number of points with the same duration and batch size stacked on top of each other.
C.4 Number of Links per Timestamp
In Figure 7, we show how often each number of links with the same timestamp (x-axis) occurs in each dataset. We can observe that all continuous-time datasets have a large number of edges that appear at a unique point in time. Additionally, the number of edges that occur at once for many continuous-time datasets is less than ten (MOOC, Wiki., Myket, Reddit) and less than 50 for most datasets except for Enron which has outliers with more than 1000 links at a certain timestamp. For discrete-time datasets, many edges occur in each snapshot.
(a) In the continuous-time temporal graphs, it is most common that at most one edge occurs per timestamp.
(b) The snapshots in discrete-time temporal graphs contain large numbers of edges, typically much larger than commonly utilized batch sizes. The Contacts dataset has fewer links per snapshot due to its much higher resolution than the remaining discrete-time datasets.
Figure 7: The link count histograms show how many edges occur per timestamp in continuous-time temporal graphs and per snapshot in discrete-time temporal graphs, respectively.
C.5 Window Sizes for Different Window Durations
Figure 8 shows the distribution of the number of links per time window for continuous- and discrete-time data sets. We can see that the distribution of window sizes depends on the distribution of edges across time as visualised in Figure 5. Datasets with a more uniform activity pattern such as Wikipedia and Reddit show a similar number of edges across windows of the same duration. Datasets with inhomogeneously distributed activities, e.g. Enron or UCI, exhibit a wide range of window sizes. For large values of $h$, i.e. horizons that include more than half of the observed time, the window sizes diverge because the window of the last time window that still includes observed links gets smaller.
(a) The number of links per window in continuous-time temporal graphs for horizons ranging from one second to one year. The $x$-axis is labelled in hours.
(b) Window sizes for discrete-time temporal graphs where $h$ represents the number of snapshots.
Figure 8: Window sizes, i.e. number of links per time window, for different horizons.
Appendix D Link Prediction using Different Batch Sizes
The results in Table 5 and 6 show the differences in AUC-ROC and average precision scores using two different batch sizes. While the same models with the same trained weights are used for both evaluations, the results show substantial differences in performance. This is in line with our expectations as described in Problem (4).
As expected due to Problem (1), the performance differences are especially large for memory-based models and the performance drops almost consistently for all datasets when using a larger batch size. Other models show no clear preference for whether larger or smaller batch sizes lead to better performance. This drop in performance for memory-based models specifically can be explained because they update their internal node representations after processing each batch, thus, larger batches lead to less frequent updates and more “out-of-date” internal representations (Su et al., 2024; Feldman & Baskin, 2024). In other words, larger batches lead to a larger information loss because more edges at the beginning of each batch are missing in the internal representation for the prediction of edges at the end of each batch.
Table 5: Mean AUC-ROC performance for dynamic link prediction ($b=800$) over five runs for the continuous- (top) and discrete-time (bottom) datasets. Values in parenthesis show the relative change compared to an evaluation using batch size $b=200$.
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
Enron 74.8( $\downarrow$3.4%) 67.9( $\downarrow$7.7%) 68.4( $\uparrow$0.5%) 59.4( $\uparrow$1.2%) 67.1( $\uparrow$1.1%) 78.5( $\downarrow$1.7%) 68.4( $\uparrow$1.1%) 82.2( $\uparrow$1.1%) 77.5( $\uparrow$1.4%) 2.1%±2.2%
UCI 79.9( $\downarrow$4.1%) 47.7( $\downarrow$7.3%) 61.0( $\downarrow$3.3%) 59.2( $\downarrow$0.7%) 57.6( $\downarrow$0.9%) 64.6( $\downarrow$6.6%) 59.6( $\downarrow$0.6%) 79.9( $\downarrow$0.9%) 76.0( $\downarrow$0.2%) 2.7%±2.7%
MOOC 82.0( $\downarrow$3.4%) 77.9( $\downarrow$3.5%) 87.5( $\downarrow$1.1%) 82.5( $\uparrow$0.2%) 70.5( $\uparrow$0.1%) 57.2( $\downarrow$7.6%) 72.6( $\downarrow$0.0%) 74.5( $\uparrow$0.1%) 81.3( $\uparrow$0.1%) 1.8%±2.6%
Wiki. 77.4( $\downarrow$5.4%) 74.5( $\downarrow$4.9%) 83.7( $\downarrow$0.5%) 83.5( $\uparrow$0.1%) 72.0( $\uparrow$0.5%) 76.5( $\downarrow$0.8%) 85.2( $\uparrow$0.1%) 88.0( $\uparrow$0.2%) 80.4( $\uparrow$0.5%) 1.4%±2.1%
LastFM 76.5( $\downarrow$2.0%) 70.4( $\downarrow$1.8%) 80.1( $\downarrow$0.8%) 68.4( $\uparrow$0.0%) 67.9( $\downarrow$0.2%) 77.4( $\downarrow$1.0%) 64.3( $\uparrow$0.1%) 65.9( $\downarrow$0.1%) 78.9( $\downarrow$0.0%) 0.7%±0.8%
Myket 64.4( $\uparrow$0.7%) 63.6( $\downarrow$1.0%) 60.7( $\downarrow$0.7%) 57.4( $\downarrow$0.3%) 32.6( $\uparrow$0.3%) 52.0( $\uparrow$0.1%) 58.4( $\downarrow$0.0%) 59.4( $\downarrow$0.2%) 33.0( $\uparrow$0.5%) 0.4%±0.3%
Social 87.3( $\downarrow$4.4%) 88.2( $\downarrow$4.8%) 92.0( $\uparrow$0.3%) 92.7( $\uparrow$0.1%) 87.8( $\uparrow$0.2%) 86.0( $\uparrow$0.2%) 95.3( $\uparrow$0.1%) 94.1( $\uparrow$0.1%) 97.4( $\uparrow$0.1%) 1.1%±2.0%
Reddit 79.0( $\downarrow$1.9%) 76.4( $\downarrow$3.9%) 80.3( $\downarrow$0.1%) 78.7( $\uparrow$0.1%) 80.3( $\uparrow$0.1%) 78.5( $\downarrow$0.2%) 76.2( $\uparrow$0.0%) 77.2( $\uparrow$0.1%) 80.2( $\uparrow$0.0%) 0.7%±1.3%
UN V. 71.3( $\downarrow$3.1%) 66.9( $\downarrow$7.8%) 67.2( $\downarrow$4.4%) 52.8( $\uparrow$0.0%) 50.2( $\uparrow$0.2%) 89.6( $\uparrow$0.1%) 53.0( $\downarrow$0.1%) 56.5( $\uparrow$0.4%) 63.3( $\uparrow$0.4%) 1.8%±2.7%
US L. 56.6( $\uparrow$0.6%) 73.5( $\downarrow$7.9%) 76.7( $\downarrow$8.7%) 78.6( $\uparrow$0.1%) 82.6( $\uparrow$0.9%) 67.2( $\downarrow$0.4%) 77.0( $\uparrow$1.9%) 90.3( $\uparrow$0.1%) 90.2( $\uparrow$0.8%) 2.4%±3.4%
UN Tr. 65.7( $\downarrow$0.6%) 63.2( $\uparrow$0.1%) 61.5( $\downarrow$2.6%) 61.4( $\downarrow$0.5%) 64.6( $\downarrow$0.2%) 86.5( $\uparrow$0.1%) 60.9( $\downarrow$0.1%) 66.2( $\downarrow$0.1%) 68.2( $\downarrow$0.2%) 0.5%±0.8%
Can. P. 62.2( $\downarrow$2.7%) 64.6( $\downarrow$3.0%) 72.9( $\downarrow$0.7%) 71.1( $\downarrow$0.7%) 67.9( $\downarrow$0.2%) 62.1( $\downarrow$1.4%) 68.2( $\uparrow$0.0%) 81.0( $\downarrow$0.2%) 93.3( $\downarrow$4.5%) 1.5%±1.6%
Flights 68.9( $\downarrow$0.9%) 68.4( $\downarrow$0.9%) 68.5( $\downarrow$0.4%) 72.7( $\uparrow$0.1%) 66.0( $\uparrow$0.0%) 74.6( $\uparrow$0.0%) 70.6( $\downarrow$0.0%) 70.7( $\downarrow$0.0%) 68.7( $\downarrow$0.0%) 0.3%±0.4%
Cont. 93.4( $\downarrow$2.2%) 93.3( $\downarrow$2.2%) 95.5( $\downarrow$0.6%) 95.5( $\uparrow$0.1%) 83.3( $\uparrow$0.1%) 91.7( $\downarrow$0.5%) 94.6( $\uparrow$0.1%) 94.2( $\uparrow$0.1%) 97.2( $\uparrow$0.1%) 0.7%±0.9%
$\mu\pm\sigma$ 2.5%±1.5% 4.1%±2.8% 1.8%±2.4% 0.3%±0.4% 0.4%±0.4% 1.5%±2.4% 0.3%±0.6% 0.3%±0.3% 0.6%±1.2%
Table 6: Mean AP performance for dynamic link prediction ($b=800$) over five runs for the continuous- (top) and discrete-time (bottom) datasets. Values in parenthesis show the relative change compared to an evaluation using batch size $b=200$.
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
Enron 68.2( $\downarrow$5.4%) 62.4( $\downarrow$10.4%) 65.2( $\downarrow$0.2%) 63.2( $\uparrow$0.1%) 66.8( $\uparrow$1.2%) 75.8( $\downarrow$1.5%) 70.7( $\uparrow$0.6%) 83.1( $\uparrow$0.9%) 77.2( $\uparrow$1.0%) 2.4%±3.4%
UCI 77.5( $\downarrow$4.5%) 46.3( $\downarrow$5.5%) 70.1( $\downarrow$1.5%) 68.3( $\downarrow$0.4%) 65.2( $\uparrow$0.1%) 61.0( $\downarrow$6.2%) 68.9( $\downarrow$0.2%) 85.5( $\downarrow$0.4%) 81.1( $\uparrow$0.6%) 2.2%±2.5%
MOOC 80.8( $\downarrow$3.1%) 74.7( $\downarrow$3.1%) 85.6( $\downarrow$1.6%) 84.6( $\uparrow$0.1%) 73.5( $\uparrow$0.0%) 56.1( $\downarrow$7.5%) 78.3( $\downarrow$0.2%) 77.8( $\downarrow$0.1%) 82.4( $\uparrow$0.0%) 1.7%±2.5%
Wiki. 79.7( $\downarrow$5.3%) 77.3( $\downarrow$4.3%) 88.5( $\downarrow$0.3%) 87.9( $\uparrow$0.1%) 75.4( $\uparrow$0.6%) 72.6( $\downarrow$0.7%) 89.6( $\uparrow$0.1%) 91.3( $\uparrow$0.2%) 83.6( $\uparrow$0.7%) 1.4%±2.0%
LastFM 75.3( $\downarrow$3.0%) 68.6( $\downarrow$3.9%) 80.1( $\downarrow$0.3%) 75.2( $\uparrow$0.3%) 69.8( $\uparrow$0.0%) 72.5( $\downarrow$1.0%) 71.9( $\uparrow$0.3%) 74.2( $\uparrow$0.2%) 81.2( $\uparrow$0.2%) 1.0%±1.4%
Myket 63.7( $\uparrow$0.4%) 61.0( $\downarrow$2.4%) 61.6( $\downarrow$0.8%) 56.6( $\downarrow$0.9%) 45.0( $\downarrow$0.3%) 51.2( $\downarrow$0.1%) 57.9( $\downarrow$0.7%) 58.8( $\downarrow$0.6%) 44.6( $\downarrow$0.3%) 0.7%±0.7%
Social 83.2( $\downarrow$5.8%) 85.7( $\downarrow$6.5%) 93.5( $\uparrow$0.3%) 94.9( $\uparrow$0.1%) 86.3( $\uparrow$0.1%) 80.8( $\uparrow$0.3%) 96.3( $\uparrow$0.1%) 95.5( $\uparrow$0.0%) 97.5( $\uparrow$0.2%) 1.5%±2.7%
Reddit 78.3( $\downarrow$2.3%) 75.7( $\downarrow$4.4%) 80.3( $\downarrow$0.3%) 78.5( $\downarrow$0.2%) 81.1( $\downarrow$0.0%) 73.5( $\downarrow$0.2%) 76.5( $\downarrow$0.1%) 77.4( $\downarrow$0.1%) 82.7( $\downarrow$0.0%) 0.8%±1.5%
UN V. 65.1( $\downarrow$4.0%) 62.2( $\downarrow$8.3%) 63.1( $\downarrow$4.5%) 51.8( $\downarrow$0.8%) 50.3( $\downarrow$0.7%) 84.9( $\uparrow$0.1%) 52.9( $\downarrow$0.7%) 53.7( $\downarrow$0.5%) 60.0( $\uparrow$0.0%) 2.2%±2.8%
US L. 48.2( $\uparrow$0.0%) 67.8( $\downarrow$7.2%) 74.4( $\downarrow$8.3%) 70.8( $\downarrow$0.6%) 81.3( $\uparrow$0.7%) 63.1( $\downarrow$0.3%) 78.8( $\uparrow$1.8%) 86.2( $\uparrow$0.2%) 86.7( $\uparrow$1.1%) 2.2%±3.2%
UN Tr. 57.9( $\downarrow$1.7%) 58.5( $\downarrow$1.2%) 57.2( $\downarrow$2.8%) 57.2( $\downarrow$1.2%) 57.4( $\downarrow$0.8%) 81.1( $\uparrow$0.1%) 55.7( $\downarrow$0.9%) 63.7( $\downarrow$0.0%) 64.5( $\uparrow$0.4%) 1.0%±0.9%
Can. P. 51.4( $\downarrow$2.6%) 60.4( $\downarrow$2.2%) 69.1( $\uparrow$0.8%) 68.1( $\uparrow$0.6%) 64.5( $\uparrow$0.9%) 61.7( $\downarrow$3.3%) 65.0( $\uparrow$1.2%) 78.9( $\uparrow$2.2%) 93.1( $\downarrow$4.1%) 2.0%±1.2%
Flights 65.8( $\downarrow$1.4%) 65.7( $\downarrow$1.5%) 68.2( $\downarrow$0.2%) 73.0( $\uparrow$0.4%) 64.8( $\downarrow$0.3%) 70.4( $\downarrow$0.1%) 70.8( $\downarrow$0.0%) 71.4( $\uparrow$0.3%) 68.4( $\downarrow$0.2%) 0.5%±0.6%
Cont. 91.4( $\downarrow$3.0%) 93.3( $\downarrow$2.3%) 95.6( $\downarrow$0.6%) 96.1( $\uparrow$0.1%) 84.5( $\uparrow$0.1%) 88.4( $\downarrow$0.5%) 95.1( $\uparrow$0.1%) 94.4( $\uparrow$0.1%) 97.8( $\uparrow$0.1%) 0.8%±1.1%
$\mu\pm\sigma$ 3.0%±1.8% 4.5%±2.8% 1.6%±2.3% 0.4%±0.4% 0.4%±0.4% 1.6%±2.4% 0.5%±0.5% 0.4%±0.6% 0.6%±1.1%
Appendix E Evaluation based on Realistic Time Horizons
In our main experiments, we determined the time horizon based on the average number of links per time window to enable a fair comparison to the batch-based approach. In Section 3.1, we discussed that the time horizon for a realistic evaluation should instead be chosen carefully for each individual dataset. To provide an example of such a realistic evaluation, we will assign a reasonable horizon for each dataset used in this work. Note that this enables a fair model comparison since we use the same horizon for all models across each dataset. The temporal resolution for most discrete-time datasets is limited due to the data collection process. The US Legislators dataset for example contains one snapshot for each congress which provides a natural horizon of one snapshot. Thus, we will only consider continuous-time datasets and discrete-time datasets where the duration of a snapshot is not yet a natural time horizon for evaluation, i.e. Contacts.
We list the considered datasets and the chosen horizon with the reasoning behind it in the following:
  • Enron is a network of users with edges representing emails sent between them. We choose 24 hours as a reasonable horizon since 90% of all email replies are typically sent within a day (Kooti et al., 2015).
  • UCI is a social network based on student communications. We use 30 min as the horizon since users receiving a text message typically feel pressured to reply between the next 20 minutes and the end of the day (Aranda & Baig, 2018).
  • MOOC connects students to units of an online course based on their interactions. We set $h=6\text{ min}$ since Guo et al. (2014) recommend keeping the learning video of a unit shorter than this time frame.
  • Wikipedia represents the editing behaviour of users in a graph. Since editing Wikipedia articles is unpaid, we do not expect frequent interactions from each user. We assume that users come from different time zones and consider that the total duration of the dataset is only one month. Therefore, we see the typical working time of 8 hours as an appropriate time horizon to take into account that interactions won’t appear very frequently but there are still enough time windows for evaluation.
  • LastFM connects users to the songs they listen to. We select 24 hours as the horizon to evaluate this dataset on a task where the goal is to predict the songs that users will listen to tomorrow based on their listening behaviour during the last days.
  • Myket represents users and Android applications that are connected when an application is installed. Considering that, similar to the Wikipedia dataset, no frequent interactions are expected, we choose 24 hours as the time horizon since the total duration of the dataset is longer than the total duration of the Wikipedia dataset.
  • Social Evolution is a proximity network gathered from students in a dormitory. Thus, we select 2 hours – a typical duration of a lecture including breaks – as the time horizon.
  • Reddit contains the posting behaviour of Reddit users. Since dynamics in a social network are typically fast, we select a time horizon of 15 minutes.
  • Contacts: Similar to Social Evolution, we select 2 hours as the time horizon since the network also captures the proximity of students.
We use the trained models from the experiments of the main part of our work and reevaluate the specified datasets with the selected time horizons. The results are presented in Table 7 and 8. For completeness, both tables also include the performance scores of the discrete-time datasets that have not been reevaluated because we determined the horizon used above as realistic.
Table 7: Average AUC-ROC performance and standard deviation over five runs using the trained models from the above experiments and the window-based evaluation with the horizons specified above. The tables also include the datasets that have not been reevaluated using performance scores obtained from the main experiments’ time horizons.
Datasets JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron 84.2 ± 4.8 80.4 ± 2.5 70.0 ± 4.5 68.3 ± 1.8 75.5 ± 0.4 83.1 ± 0.0 74.2 ± 4.8 87.9 ± 0.6 83.9 ± 0.6
UCI 89.4 ± 0.7 70.4 ± 2.8 64.0 ± 1.3 52.2 ± 1.1 56.0 ± 0.2 75.3 ± 0.0 55.3 ± 1.3 82.7 ± 0.9 75.4 ± 0.4
MOOC 84.6 ± 4.1 81.3 ± 3.3 87.6 ± 1.3 80.7 ± 0.9 69.3 ± 1.5 62.9 ± 0.0 68.1 ± 0.9 70.4 ± 1.6 79.9 ± 10.0
Wiki. 75.5 ± 0.8 73.2 ± 0.9 83.5 ± 0.6 83.5 ± 0.2 71.9 ± 0.8 76.1 ± 0.0 85.2 ± 0.6 87.9 ± 0.3 80.6 ± 1.6
LastFM 74.9 ± 2.1 67.5 ± 1.4 78.6 ± 2.9 66.0 ± 0.9 67.2 ± 0.3 77.5 ± 0.0 63.2 ± 6.7 60.6 ± 1.3 78.9 ± 0.6
Myket 64.5 ± 1.8 62.7 ± 3.1 60.4 ± 2.3 57.8 ± 0.4 32.2 ± 0.4 51.4 ± 0.0 58.3 ± 2.0 59.9 ± 0.3 32.8 ± 0.9
Social 87.0 ± 2.0 84.4 ± 4.7 92.3 ± 2.7 92.6 ± 0.5 86.7 ± 0.0 85.1 ± 0.0 94.7 ± 0.5 94.7 ± 0.2 97.4 ± 0.1
Reddit 80.6 ± 0.1 79.5 ± 0.8 80.4 ± 0.4 78.6 ± 0.7 80.2 ± 0.3 78.6 ± 0.0 76.2 ± 0.4 77.1 ± 0.4 80.2 ± 1.1
Avg. Rank 3.0 5.2 3.8 6.1 6.9 6.0 6.1 4.2 3.6
UN V. 54.0 ± 1.8 52.2 ± 2.0 51.3 ± 7.1 54.4 ± 3.6 53.7 ± 2.1 89.6 ± 0.0 53.4 ± 1.0 56.9 ± 1.6 65.2 ± 1.1
US L. 52.5 ± 1.8 61.8 ± 3.5 57.7 ± 1.8 78.6 ± 7.9 82.0 ± 4.0 68.4 ± 0.0 75.4 ± 5.3 90.4 ± 1.5 89.4 ± 0.9
UN Tr. 57.7 ± 3.3 50.3 ± 1.4 54.3 ± 1.5 64.0 ± 1.3 67.6 ± 1.2 85.6 ± 0.0 63.7 ± 1.6 68.6 ± 2.6 70.7 ± 2.6
Can. P. 63.6 ± 0.8 67.5 ± 8.5 73.2 ± 1.1 72.7 ± 2.2 70.0 ± 1.4 63.2 ± 0.0 69.5 ± 3.1 80.7 ± 0.9 85.5 ± 3.5
Flights 67.4 ± 2.0 66.0 ± 1.9 68.1 ± 1.7 72.6 ± 0.2 66.0 ± 1.7 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.8 ± 1.0
Cont. 85.4 ± 0.5 74.5 ± 3.1 94.6 ± 0.6 96.0 ± 0.2 86.6 ± 0.1 85.8 ± 0.0 95.7 ± 0.4 95.2 ± 0.2 97.8 ± 0.0
Avg. Rank 7.3 8.0 6.5 3.5 5.5 4.2 5.2 2.7 2.2
The results using both the AUC-ROC as well as the average precision score mostly agree on a best-performing model, yet for different datasets, there is no clear winner among the models. For continuous-time datasets, JODIE, GraphMixer, and DyGFormer are among the best-performing models while EdgeBank, GraphMixer, and DyGFormer performed best for discrete-time data.
DyGFormer performs best for both Contacts and Social Evolution, suggesting that DyGFormer is best suited for proximity networks among all models. For UN Vote, UN Trade, and Flights, EdgeBank – a simple baseline model that predicts an edge if it has occurred before – is among the best. These datasets are highly repetitive because they are, e.g. based on a schedule or relations among countries that rarely change. Thus, since they are all outperformed by simple baselines, none of the proposed TGNN models adequately address the task of these datasets, i.e. finding edges that do not follow the schedule or some other reoccurring pattern. For other types of datasets like communication (e.g. Enron or UCI) or user-interaction networks (e.g. Wikipedia or MOOC), no clear patterns are visible.
Table 8: Mean average precision scores and standard deviation following Table 7.
Datasets JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron 82.2 ± 4.8 80.0 ± 3.4 71.2 ± 4.2 72.6 ± 1.3 77.1 ± 0.3 81.6 ± 0.0 77.9 ± 2.6 89.2 ± 0.4 85.0 ± 0.6
UCI 93.0 ± 0.5 81.5 ± 1.9 78.0 ± 0.9 71.6 ± 0.7 73.3 ± 0.2 75.6 ± 0.0 73.2 ± 0.7 89.6 ± 0.5 85.2 ± 0.3
MOOC 84.1 ± 5.4 79.4 ± 3.8 87.4 ± 1.6 83.9 ± 0.8 73.8 ± 1.1 62.1 ± 0.0 75.8 ± 0.5 75.6 ± 0.9 82.6 ± 8.9
Wiki. 78.3 ± 1.2 76.4 ± 0.7 88.3 ± 0.4 88.0 ± 0.2 75.7 ± 1.1 72.4 ± 0.0 89.7 ± 0.4 91.3 ± 0.2 84.0 ± 1.2
LastFM 73.6 ± 2.0 64.8 ± 1.8 78.6 ± 3.7 73.0 ± 0.8 69.2 ± 0.5 72.9 ± 0.0 71.2 ± 6.7 71.0 ± 1.1 81.3 ± 0.9
Myket 62.9 ± 1.5 60.2 ± 1.9 61.1 ± 1.7 56.8 ± 0.4 44.9 ± 0.2 51.1 ± 0.0 57.8 ± 2.2 59.0 ± 0.2 44.4 ± 1.7
Social 82.5 ± 4.0 81.6 ± 5.7 94.0 ± 1.8 95.0 ± 0.3 85.8 ± 0.1 79.9 ± 0.0 96.2 ± 0.3 95.8 ± 0.2 97.8 ± 0.1
Reddit 80.1 ± 0.3 79.2 ± 0.9 80.6 ± 0.6 78.6 ± 1.0 81.3 ± 0.4 73.5 ± 0.0 76.5 ± 0.6 77.5 ± 0.5 82.8 ± 0.8
Avg. Rank 3.4 5.8 3.8 5.5 6.8 7.2 5.4 4.0 3.2
UN V. 52.6 ± 1.8 49.6 ± 1.9 49.7 ± 3.9 52.7 ± 2.6 52.4 ± 2.0 84.2 ± 0.0 52.4 ± 0.9 54.0 ± 1.4 62.4 ± 1.7
US L. 46.0 ± 0.9 62.5 ± 3.6 58.6 ± 2.4 71.0 ± 8.9 80.7 ± 3.7 63.2 ± 0.0 77.5 ± 4.3 86.5 ± 1.9 86.1 ± 1.0
UN Tr. 52.7 ± 3.0 49.4 ± 0.9 53.2 ± 1.5 58.9 ± 2.7 59.2 ± 1.7 79.0 ± 0.0 57.5 ± 1.9 65.8 ± 1.9 67.1 ± 2.7
Can. P. 52.1 ± 0.5 61.0 ± 7.6 69.9 ± 0.8 70.8 ± 1.6 68.3 ± 2.3 59.4 ± 0.0 68.2 ± 1.6 80.9 ± 0.5 83.2 ± 2.9
Flights 65.2 ± 2.7 63.9 ± 2.8 68.3 ± 2.2 73.5 ± 0.3 65.5 ± 1.3 70.4 ± 0.0 71.0 ± 0.4 71.9 ± 0.8 69.1 ± 1.8
Cont. 83.3 ± 0.7 69.7 ± 3.7 95.0 ± 0.8 96.9 ± 0.2 88.1 ± 0.2 82.3 ± 0.0 96.7 ± 0.5 95.8 ± 0.1 98.5 ± 0.0
Avg. Rank 7.7 8.3 6.3 3.3 5.2 4.7 4.8 2.5 2.2
Appendix F Experimental Details
For reproducibility, we provide a GitHub repository [^3] https://github.com/M-Lampert/DyGLib extending the dynamic graph learning library DyGLib [^4] https://github.com/yule-BUAA/DyGLib (MIT License) (Yu et al., 2023), including a bash script to run the experiments.
We use the best hyperparameters reported by Yu et al. (2023) and, for completeness, list these hyperparameters for the 13 datasets used by Yu et al. (2023) below. However, the Myket dataset (Loghmani & Fazli, 2023) was not included in the study. Therefore, for Myket, we use each method’s default parameters as suggested by the respective authors.
We use 9 state-of-the-art dynamic graph learning models and baselines (JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), CAWN (Wang et al., 2021b), EdgeBank (Poursafaei et al., 2022), TCL (Wang et al., 2021a), GraphMixer (Cong et al., 2023) and DyGFormer (Yu et al., 2023)). The neural-network-based approaches (all except EdgeBank) are trained five times for 100 epochs using the Adam optimizer with a learning rate of $0.0001$. An early-stopping strategy with a patience of 5 is employed to avoid overfitting. For training and validation, a batch size of 200 is used. The training, validation and test sets of each dataset contain 70%, 15% and 15% of the edges, respectively. The sets are split based on time, i.e., the training set contains the edges that occurred first while the test set comprises the most recent edges.
The experiments were conducted on a variety of machines with different CPUs and GPUs. A list of machine specifications is provided in Table 9.
Table 9: Hardware details of the machines used for the experiments.
(a) CPUs
CPU
AMD Ryzen Threadripper PRO 5965WX 24 Cores
AMD Ryzen 9 7900X 12 Cores
11th Gen Intel(R) Core(TM) i9-11900K 8 Cores
AMD Ryzen 9 7950X 16 Cores
13th Gen Intel(R) Core(TM) i9-13900H 14 Cores
AMD Epyc 7543 (3rd Gen)
AMD Epyc 7763 (3rd Gen)
AMD Epyc 7713 (3rd Gen)
Intel Xeon Platinum 8480+ (4th Gen)
(b) GPUs
GPU
NVIDIA GeForce RTX 3090 Ti
NVIDIA GeForce RTX 4080
NVIDIA GeForce RTX 3090
NVIDIA GeForce RTX 4090
NVIDIA GeForce RTX 4060 (Laptop)
NVIDIA A100
NVIDIA GeForce RTX 2080 Ti
NVIDIA TITAN Xp
NVIDIA TITAN X
NVIDIA Quadro RTX 8000
NVIDIA H100
NVIDIA L40
NVIDIA L40s
For all model architectures, time-related representations use a size of 100 dimensions while all other non-time-related representations are set to 172. An exception is DyGFormer where the neighbor co-occurrence encoding and the aligned encoding each have 50 dimensions. We use eight attention heads for CAWN, and two attention heads for all other attention-based methods. The memory-based models either use a vanilla recurrent neural network (JODIE and DyRep), or a gated recurrent unit (GRU) to update their memory. Other model-specific parameters are provided in Table 10
Table 10: Specific hyperparameters for different models and datasets.
(a) Hyperparameters for neighborhood sampling-based models. $n_{\text{Neighbors}}$ is the number of sampled neighbors using the specified neighbor sampling strategy. $n_{\text{Layers}}$ is the number of transformer layers (for TCL), the number of MLP-Mixer layers (for GraphMixer) or the number of GNN layers otherwise.
Dataset Model Neigh. Sampling $n_{\text{Neighbors}}$ $n_{\text{Layers}}$ Dropout
Wikipedia DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 30 2 0.500 $0$.5000000.500\,$0$.500 000
Reddit DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT uniform 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL uniform 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 10 2 0.500 $0$.5000000.500\,$0$.500 000
MOOC DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.200 $0$.2000000.200\,$0$.200 000
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 20 2 0.400 $0$.4000000.400\,$0$.400 000
LastFM DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.300 $0$.3000000.300\,$0$.300 000
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 10 2 0.000 $00\,0000$
Enron DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.200 $0$.2000000.200\,$0$.200 000
TGN recent 10 1 0.000 $00\,0000$
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 20 2 0.500 $0$.5000000.500\,$0$.500 000
Social Evo. DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.000 $00\,0000$
TCL recent 20 2 0.000 $00\,0000$
GraphMixer recent 20 2 0.300 $0$.3000000.300\,$0$.300 000
UCI DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL recent 20 2 0.000 $00\,0000$
GraphMixer recent 20 2 0.400 $0$.4000000.400\,$0$.400 000
Myket DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
Flights DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
GraphMixer recent 20 2 0.200 $0$.2000000.200\,$0$.200 000
Can. Parl. DyRep uniform 10 1 0.000 $00\,0000$
TGAT uniform 20 2 0.200 $0$.2000000.200\,$0$.200 000
TGN uniform 10 1 0.300 $0$.3000000.300\,$0$.300 000
TCL uniform 20 2 0.200 $0$.2000000.200\,$0$.200 000
GraphMixer uniform 20 2 0.200 $0$.2000000.200\,$0$.200 000
US Legis. DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL uniform 20 2 0.300 $0$.3000000.300\,$0$.300 000
GraphMixer recent 20 2 0.400 $0$.4000000.400\,$0$.400 000
UN Trade DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT uniform 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.200 $0$.2000000.200\,$0$.200 000
TCL uniform 20 2 0.000 $00\,0000$
GraphMixer uniform 20 2 0.100 $0$.1000000.100\,$0$.100 000
UN Vote DyRep recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TGAT recent 20 2 0.200 $0$.2000000.200\,$0$.200 000
TGN uniform 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL uniform 20 2 0.000 $00\,0000$
GraphMixer uniform 20 2 0.000 $00\,0000$
Contacts DyRep recent 10 1 0.000 $00\,0000$
TGAT recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
TGN recent 10 1 0.100 $0$.1000000.100\,$0$.100 000
TCL recent 20 2 0.000 $00\,0000$
GraphMixer recent 20 2 0.100 $0$.1000000.100\,$0$.100 000
(b) Hyperparameters DyGFormer.
Dataset Model Sequence Length Patch Size Dropout
Wikipedia DyGFormer 32 1 0.100 $0$.1000000.100\,$0$.100 000
Reddit DyGFormer 64 2 0.200 $0$.2000000.200\,$0$.200 000
MOOC DyGFormer 256 8 0.100 $0$.1000000.100\,$0$.100 000
LastFM DyGFormer 512 16 0.100 $0$.1000000.100\,$0$.100 000
Enron DyGFormer 256 8 0.000 $00\,0000$
Social Evo. DyGFormer 32 1 0.100 $0$.1000000.100\,$0$.100 000
UCI DyGFormer 32 1 0.100 $0$.1000000.100\,$0$.100 000
Myket DyGFormer 32 1 0.100 $0$.1000000.100\,$0$.100 000
Flights DyGFormer 256 8 0.100 $0$.1000000.100\,$0$.100 000
Can. Parl. DyGFormer 2048 64 0.100 $0$.1000000.100\,$0$.100 000
US Legis. DyGFormer 256 8 0.000 $00\,0000$
UN Trade DyGFormer 256 8 0.000 $00\,0000$
UN Vote DyGFormer 128 4 0.200 $0$.2000000.200\,$0$.200 000
Contacts DyGFormer 32 1 0.000 $00\,0000$
(c) Hyperparameters CAWN.
Dataset Model Walk Length Time Scale Dropout
Wikipedia CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Reddit CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
MOOC CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
LastFM CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Enron CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Social Evo. CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
UCI CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Myket CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Flights CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Can. Parl. CAWN 1 0.000001 0.000 $00\,0000$
US Legis. CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
UN Trade CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
UN Vote CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
Contacts CAWN 1 0.000001 0.100 $0$.1000000.100\,$0$.100 000
(d) Hyperparameters EdgeBank
Dataset Model Neg. Sampling Memory Mode Time Window
Wikipedia EdgeBank random unlimited -
historical repeat threshold -
inductive repeat threshold -
Reddit EdgeBank random unlimited -
historical repeat threshold -
inductive repeat threshold -
MOOC EdgeBank random time window fixed proportion
historical time window repeat interval
inductive repeat threshold -
LastFM EdgeBank random time window fixed proportion
historical time window repeat interval
inductive repeat threshold -
Enron EdgeBank random time window fixed proportion
historical time window repeat interval
inductive repeat threshold -
Social Evo. EdgeBank random repeat threshold -
historical repeat threshold -
inductive repeat threshold -
UCI EdgeBank random unlimited -
historical time window fixed proportion
inductive time window repeat interval
Myket EdgeBank random unlimited -
historical repeat threshold
inductive repeat threshold
Flights EdgeBank random unlimited -
historical repeat threshold -
inductive repeat threshold -
Can. Parl. EdgeBank random time window fixed proportion
historical time window fixed proportion
inductive repeat threshold -
US Legis. EdgeBank random time window fixed proportion
historical time window fixed proportion
inductive time window fixed proportion
UN Trade EdgeBank random time window repeat interval
historical time window repeat interval
inductive repeat threshold -
UN Vote EdgeBank random time window repeat interval
historical time window repeat interval
inductive time window repeat interval
Contacts EdgeBank random time window repeat interval
historical time window repeat interval
inductive repeat threshold -
Appendix G Detailed AUC-ROC and Average Precision Results
Here, we provide the evaluation runtime of all models for dynamic link prediction and forecasting averaged across five runs and detailed tabulated results for all models’ AUC-ROC and average precision performance across five runs, including standard deviations. We further include visualisations investigating the causes of the large performance differences between dynamic link prediction and forecasting.
G.1 Runtime
Table 11 shows the runtime during the evaluation of all models for dynamic link prediction and forecasting. The runtime of both approaches is generally comparable with both approaches being faster than the other on some datasets. For continuous-time datasets, we can see that the runtime is often comparable and large differences are caused by specific characteristics of the datasets. The model evaluation on UCI, for example, is faster using the batch-based approach because it consistently evaluates batches of size $b=200$ while the time windows are mostly smaller which leads to less parallelization. This is because the duration of each window was chosen so that the average window size across the whole dataset is comparable to the chosen batch size of $b=200$. For this specific dataset, there are large bursts of activity in the first third of the total dataset duration and low activity afterwards. This leads to small window sizes for the temporal edges of the test set (see Figure 5). For discrete-time datasets, each snapshot contains more than 200 edges for most datasets leading to better parallelization for link forecasting and, thus, smaller runtime.
Table 11: Mean and standard deviation in seconds of the evaluation runtime of all models for dynamic link prediction and forecasting averaged across five runs.
Dataset Approach JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron Forec. 3.42 ± 0.26 4.15 ± 0.30 4.20 ± 0.31 36.16 ± 0.35 67.98 ± 0.76 9.00 ± 0.11 7.76 ± 0.26 12.44 ± 0.33 43.58 ± 0.43
Pred. 3.47 ± 1.57 4.16 ± 1.99 4.29 ± 1.92 37.26 ± 2.03 73.06 ± 1.38 7.74 ± 0.09 7.61 ± 1.18 12.65 ± 1.24 47.79 ± 1.69
UCI Forec. 6.84 ± 0.47 8.40 ± 0.50 8.44 ± 0.50 37.72 ± 0.44 97.97 ± 1.58 4.56 ± 0.17 11.36 ± 0.30 13.90 ± 0.31 23.85 ± 0.39
Pred. 2.66 ± 1.02 3.08 ± 0.72 3.09 ± 0.72 24.58 ± 0.83 88.00 ± 2.72 1.15 ± 0.02 4.86 ± 0.97 8.03 ± 0.99 18.29 ± 0.78
MOOC Forec. 55.41 ± 0.52 59.23 ± 0.37 58.56 ± 0.38 210.16 ± 0.28 539.03 ± 2.59 264.74 ± 2.79 89.27 ± 0.20 132.41 ± 0.42 226.80 ± 1.38
Pred. 61.10 ± 0.56 64.53 ± 0.43 65.49 ± 0.35 230.79 ± 1.74 555.57 ± 3.37 301.55 ± 5.18 95.60 ± 1.19 140.46 ± 1.36 254.47 ± 0.93
Wiki. Forec. 15.16 ± 0.73 16.29 ± 0.79 16.46 ± 0.65 44.58 ± 0.88 90.17 ± 1.10 10.27 ± 0.46 17.19 ± 0.70 19.80 ± 0.71 50.72 ± 1.13
Pred. 13.67 ± 1.08 14.63 ± 0.84 13.91 ± 0.91 45.38 ± 1.74 89.25 ± 1.72 9.46 ± 0.43 15.99 ± 1.44 19.01 ± 1.46 49.27 ± 1.50
LastFM Forec. 271.34 ± 0.97 281.58 ± 1.81 338.74 ± 1.73 985.54 ± 34.93 3316.72 ± 21.18 $1$ 18.08 578.48 ± 3.48 702.40 ± 1.33 1037.08 ± 8.49
Pred. 319.89 ± 1.76 340.54 ± 12.63 400.65 ± 11.46 1180.93 ± 83.27 3472.23 ± 79.24 1723.99 ± 28.39 674.63 ± 12.96 832.52 ± 3.00 1193.28 ± 3.40
Myket Forec. 467.24 ± 3.94 465.55 ± 6.40 445.55 ± 11.19 680.26 ± 15.71 851.62 ± 21.51 479.92 ± 6.43 507.72 ± 7.10 602.63 ± 7.31 643.33 ± 8.06
Pred. 425.36 ± 6.72 424.50 ± 7.38 403.12 ± 9.27 701.21 ± 8.81 876.41 ± 17.09 445.81 ± 5.62 475.26 ± 8.32 566.69 ± 19.39 628.24 ± 7.03
Social Forec. 408.85 ± 1.89 424.96 ± 3.62 419.62 ± 2.44 1208.04 ± 12.30 4020.21 ± 4.69 1090.46 ± 1.08 709.80 ± 2.45 899.58 ± 2.20 1025.54 ± 2.26
Pred. 500.15 ± 3.65 509.14 ± 2.35 510.64 ± 3.11 1544.19 ± 2.59 4308.14 ± 8.00 1379.71 ± 0.74 955.07 ± 1.34 1203.01 ± 1.67 1263.80 ± 4.42
Reddit Forec. 120.79 ± 0.77 124.65 ± 0.74 131.35 ± 2.35 543.41 ± 1.25 561.97 ± 2.74 174.94 ± 0.50 161.26 ± 1.12 213.45 ± 0.73 248.81 ± 1.26
Pred. 120.22 ± 0.46 124.28 ± 0.41 129.58 ± 0.67 585.98 ± 53.62 575.36 ± 2.94 191.38 ± 0.59 167.83 ± 1.06 227.86 ± 2.07 254.34 ± 1.45
UN V. Forec. 4.25 ± 0.30 9.37 ± 0.28 20.11 ± 0.36 161.34 ± 1.82 5804.24 ± 9.62 6.88 ± 0.05 40.01 ± 0.29 831.89 ± 9.65 137.97 ± 0.97
Pred. 158.35 ± 0.46 161.49 ± 0.70 174.23 ± 0.68 560.17 ± 4.39 5599.27 ± 92.68 684.81 ± 8.57 310.53 ± 0.78 1223.42 ± 8.09 406.54 ± 0.94
US L. Forec. 1.01 ± 0.26 1.33 ± 0.31 1.32 ± 0.31 13.25 ± 0.30 32.43 ± 0.55 0.05 ± 0.01 3.19 ± 0.25 3.63 ± 0.26 18.67 ± 0.26
Pred. 2.55 ± 1.82 2.93 ± 1.76 2.84 ± 1.77 19.40 ± 0.82 42.72 ± 1.06 0.71 ± 0.02 5.73 ± 0.98 6.87 ± 1.25 26.85 ± 1.21
UN Tr. Forec. 1.95 ± 0.26 4.13 ± 0.24 3.97 ± 0.12 271.15 ± 3.19 2813.05 ± 2.68 1.83 ± 0.06 19.42 ± 0.26 507.83 ± 14.74 84.71 ± 0.52
Pred. 34.63 ± 0.42 37.36 ± 0.44 36.37 ± 0.44 466.44 ± 2.36 2992.96 ± 384.16 161.85 ± 0.63 90.81 ± 0.43 719.00 ± 70.07 213.38 ± 1.02
Can. P. Forec. 0.56 ± 0.22 2.00 ± 0.24 1.99 ± 0.24 60.15 ± 0.31 1034.22 ± 1.06 0.11 ± 0.02 4.11 ± 0.21 20.65 ± 0.30 464.45 ± 25.61
Pred. 2.38 ± 0.46 3.68 ± 0.54 3.80 ± 0.56 90.50 ± 1.56 1242.03 ± 185.41 2.02 ± 0.03 7.64 ± 0.63 30.36 ± 1.72 380.33 ± 2.02
Flights Forec. 343.75 ± 10.64 346.74 ± 9.28 341.56 ± 9.66 357.77 ± 5.58 10363.33 ± 2098.15 21.36 ± 0.44 68.14 ± 0.78 1611.70 ± 33.57 994.11 ± 7.16
Pred. 1567.52 ± 79.75 1550.22 ± 94.86 1581.06 ± 47.29 2359.88 ± 54.11 3420.00 ± 57.27 2132.99 ± 103.26 1933.46 ± 84.54 2113.83 ± 20.74 2701.10 ± 31.46
Cont. Forec. 656.31 ± 2.85 673.36 ± 3.62 683.82 ± 2.96 1966.25 ± 10.58 3878.32 ± 11.95 2646.55 ± 37.98 1139.57 ± 4.54 1378.58 ± 4.22 1518.44 ± 12.00
Pred. 938.48 ± 6.81 940.55 ± 7.94 953.46 ± 8.02 2528.31 ± 34.70 4358.78 ± 21.10 3854.40 ± 32.23 1523.29 ± 26.55 1848.28 ± 40.53 1873.09 ± 11.28
G.2 AUC-ROC
The following reports the complete results (including standard deviations) of the experiments conducted in Section 4.
Table 12: Average AUC-ROC performance over five runs for the test set of the continuous-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. Enron 84.0 ± 5.1 80.3 ± 1.4 67.9 ± 7.1 69.0 ± 1.6 75.7 ± 0.5 82.7 ± 0.0 75.1 ± 5.2 88.6 ± 0.5 84.5 ± 0.6
UCI 86.8 ± 1.0 60.2 ± 2.8 62.1 ± 1.3 55.2 ± 1.4 56.5 ± 0.5 72.5 ± 0.0 56.3 ± 1.0 80.2 ± 1.0 75.7 ± 0.5
MOOC 83.1 ± 4.2 79.0 ± 4.5 87.4 ± 1.9 79.9 ± 0.8 68.8 ± 1.6 59.8 ± 0.0 68.4 ± 1.4 70.3 ± 1.2 80.0 ± 9.0
Wiki. 81.5 ± 0.4 78.3 ± 0.4 83.7 ± 0.6 82.9 ± 0.3 71.3 ± 0.8 77.2 ± 0.0 84.6 ± 0.5 87.3 ± 0.3 79.8 ± 1.6
LastFM 76.3 ± 0.8 69.0 ± 1.4 79.2 ± 2.7 65.2 ± 0.9 66.3 ± 0.3 78.0 ± 0.0 62.5 ± 6.4 59.9 ± 1.4 78.2 ± 0.6
Myket 64.4 ± 2.2 64.1 ± 2.9 61.2 ± 2.6 57.8 ± 0.5 33.5 ± 0.4 52.6 ± 0.0 58.3 ± 2.2 59.8 ± 0.4 33.8 ± 0.9
Social 92.1 ± 1.9 92.2 ± 0.7 92.2 ± 2.6 92.5 ± 0.5 86.5 ± 0.0 84.9 ± 0.0 94.7 ± 0.5 94.6 ± 0.2 97.3 ± 0.1
Reddit 80.6 ± 0.1 79.5 ± 0.8 80.4 ± 0.4 78.6 ± 0.7 80.2 ± 0.3 78.6 ± 0.0 76.2 ± 0.4 77.1 ± 0.4 80.2 ± 1.1
Avg. Rank 3.0 5.1 3.6 6.0 7.0 6.4 6.1 4.2 3.5
Pred. Enron 77.4 ± 3.6 73.5 ± 2.4 68.0 ± 2.9 58.7 ± 1.2 66.4 ± 0.4 79.8 ± 0.0 67.6 ± 5.5 81.3 ± 0.8 76.4 ± 0.5
UCI 83.3 ± 1.4 51.4 ± 7.8 63.0 ± 1.3 59.6 ± 1.5 58.2 ± 0.6 69.1 ± 0.0 60.0 ± 0.9 80.6 ± 0.8 76.2 ± 0.6
MOOC 84.8 ± 3.1 80.7 ± 3.2 88.5 ± 1.6 82.3 ± 0.6 70.4 ± 1.3 61.9 ± 0.0 72.6 ± 0.6 74.4 ± 1.4 81.2 ± 8.9
Wiki. 81.8 ± 0.4 78.4 ± 0.4 84.1 ± 0.6 83.5 ± 0.2 71.6 ± 0.8 77.1 ± 0.0 85.2 ± 0.5 87.8 ± 0.3 80.0 ± 1.6
LastFM 78.0 ± 0.7 71.7 ± 1.1 80.7 ± 2.4 68.4 ± 0.7 68.1 ± 0.3 78.2 ± 0.0 64.3 ± 6.0 65.9 ± 1.7 78.9 ± 0.6
Myket 64.0 ± 2.1 64.2 ± 2.7 61.1 ± 2.6 57.6 ± 0.4 32.5 ± 0.4 52.0 ± 0.0 58.4 ± 2.0 59.5 ± 0.4 32.8 ± 1.0
Social 91.3 ± 2.1 92.7 ± 0.5 91.7 ± 3.3 92.6 ± 0.5 87.7 ± 0.1 85.8 ± 0.0 95.2 ± 0.2 94.1 ± 0.2 97.3 ± 0.1
Reddit 80.6 ± 0.1 79.5 ± 0.8 80.4 ± 0.4 78.7 ± 0.6 80.2 ± 0.3 78.6 ± 0.0 76.2 ± 0.4 77.1 ± 0.4 80.2 ± 1.1
Avg. Rank 3.1 5.1 3.4 5.8 7.6 6.1 5.9 4.1 3.9
Table 13: Average AUC-ROC performance over five runs for the test set of the discrete-time datasets, including standard deviation. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. UN V. 54.0 ± 1.8 52.2 ± 2.0 51.3 ± 7.1 54.4 ± 3.6 53.7 ± 2.1 89.6 ± 0.0 53.4 ± 1.0 56.9 ± 1.6 65.2 ± 1.1
US L. 52.5 ± 1.8 61.8 ± 3.5 57.7 ± 1.8 78.6 ± 7.9 82.0 ± 4.0 68.4 ± 0.0 75.4 ± 5.3 90.4 ± 1.5 89.4 ± 0.9
UN Tr. 57.7 ± 3.3 50.3 ± 1.4 54.3 ± 1.5 64.0 ± 1.3 67.6 ± 1.2 85.6 ± 0.0 63.7 ± 1.6 68.6 ± 2.6 70.7 ± 2.6
Can. P. 63.6 ± 0.8 67.5 ± 8.5 73.2 ± 1.1 72.7 ± 2.2 70.0 ± 1.4 63.2 ± 0.0 69.5 ± 3.1 80.7 ± 0.9 85.5 ± 3.5
Flights 67.4 ± 2.0 66.0 ± 1.9 68.1 ± 1.7 72.6 ± 0.2 66.0 ± 1.7 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.8 ± 1.0
Cont. 95.6 ± 0.8 94.9 ± 0.3 96.6 ± 0.3 95.9 ± 0.2 86.7 ± 0.1 93.0 ± 0.0 95.7 ± 0.5 95.2 ± 0.2 97.7 ± 0.0
Avg. Rank 6.8 7.7 6.0 3.7 6.0 4.3 5.3 3.0 2.2
Pred. UN V. 73.7 ± 2.4 72.6 ± 1.5 70.3 ± 4.3 52.8 ± 3.6 50.1 ± 1.6 89.5 ± 0.0 53.0 ± 1.6 56.2 ± 2.0 63.0 ± 1.1
US L. 56.3 ± 1.9 79.9 ± 1.1 84.0 ± 2.2 78.5 ± 7.8 81.8 ± 4.0 67.5 ± 0.0 75.6 ± 5.4 90.2 ± 1.6 89.4 ± 0.9
UN Tr. 66.1 ± 3.0 63.2 ± 2.1 63.1 ± 1.2 61.7 ± 1.3 64.7 ± 1.3 86.4 ± 0.0 60.9 ± 1.3 66.3 ± 2.5 68.3 ± 2.3
Can. P. 63.9 ± 0.7 66.6 ± 2.5 73.4 ± 3.5 71.6 ± 2.6 68.0 ± 1.0 62.9 ± 0.0 68.2 ± 3.6 81.2 ± 1.0 97.7 ± 0.7
Flights 69.5 ± 2.2 69.0 ± 1.0 68.8 ± 1.6 72.6 ± 0.2 66.0 ± 1.7 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.8 ± 1.0
Cont. 95.5 ± 0.5 95.4 ± 0.2 96.1 ± 0.8 95.4 ± 0.3 83.3 ± 0.0 92.2 ± 0.0 94.5 ± 0.5 94.1 ± 0.2 97.1 ± 0.0
Avg. Rank 5.2 5.3 4.3 5.3 7.0 4.7 6.3 3.7 3.2
G.3 Average precision
Average precision scores as an additional metric to evaluate experiments in Section 4. The results are comparable to the results using the AUC-ROC score.
Table 14: Mean average precision performance for dynamic link forecasting (window-based) over five runs for the continuous-time datasets. Values in parenthesis show the relative change as compared to the average precision performance for dynamic link prediction (batch-based).
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
Enron 80.8( $\uparrow$12.0%) 78.3( $\uparrow$12.4%) 68.6( $\uparrow$5.0%) 71.7( $\uparrow$13.5%) 76.0( $\uparrow$15.1%) 81.1( $\uparrow$5.5%) 78.2( $\uparrow$11.3%) 89.8( $\uparrow$9.2%) 85.3( $\uparrow$11.6%) 10.6%±3.4%
UCI 87.0( $\uparrow$7.2%) 59.5( $\uparrow$21.4%) 69.2( $\downarrow$2.8%) 64.4( $\downarrow$6.2%) 64.0( $\downarrow$1.6%) 68.6( $\uparrow$5.4%) 65.2( $\downarrow$5.5%) 85.3( $\downarrow$0.6%) 80.5( $\downarrow$0.2%) 5.7%±6.4%
MOOC 82.4( $\downarrow$1.1%) 76.9( $\downarrow$0.3%) 86.3( $\downarrow$0.8%) 82.7( $\downarrow$2.1%) 72.3( $\downarrow$1.7%) 59.0( $\downarrow$2.7%) 74.9( $\downarrow$4.6%) 74.4( $\downarrow$4.5%) 82.1( $\downarrow$0.4%) 2.0%±1.6%
Wiki. 84.1( $\downarrow$0.1%) 80.9( $\uparrow$0.1%) 88.5( $\downarrow$0.4%) 87.5( $\downarrow$0.4%) 75.1( $\uparrow$0.1%) 73.3( $\uparrow$0.3%) 89.2( $\downarrow$0.4%) 90.8( $\downarrow$0.4%) 83.1( $\uparrow$0.0%) 0.2%±0.2%
LastFM 76.7( $\downarrow$1.1%) 69.4( $\downarrow$2.7%) 78.8( $\downarrow$1.9%) 72.1( $\downarrow$3.9%) 68.2( $\downarrow$2.3%) 73.4( $\uparrow$0.3%) 70.3( $\downarrow$1.8%) 70.0( $\downarrow$5.4%) 80.5( $\downarrow$0.7%) 2.2%±1.6%
Myket 64.5( $\uparrow$1.6%) 63.1( $\uparrow$0.9%) 62.8( $\uparrow$1.2%) 57.9( $\uparrow$1.4%) 46.6( $\uparrow$3.3%) 51.9( $\uparrow$1.3%) 58.9( $\uparrow$1.0%) 60.0( $\uparrow$1.5%) 46.1( $\uparrow$3.3%) 1.7%±0.9%
Social 89.4( $\uparrow$1.3%) 91.9( $\uparrow$0.3%) 93.9( $\uparrow$0.8%) 95.0( $\uparrow$0.2%) 85.6( $\downarrow$0.6%) 79.7( $\downarrow$1.1%) 96.1( $\downarrow$0.1%) 95.8( $\uparrow$0.4%) 97.7( $\uparrow$0.3%) 0.6%±0.4%
Reddit 80.1( $\downarrow$0.0%) 79.2( $\uparrow$0.0%) 80.6( $\uparrow$0.0%) 78.6( $\downarrow$0.1%) 81.3( $\uparrow$0.2%) 73.5( $\downarrow$0.2%) 76.5( $\downarrow$0.0%) 77.5( $\downarrow$0.0%) 82.8( $\uparrow$0.1%) 0.1%±0.1%
$\mu\pm\sigma$ 3.0%±4.3% 4.8%±7.9% 1.6%±1.6% 3.5%±4.6% 3.1%±5.0% 2.1%±2.2% 3.1%±3.9% 2.8%±3.3% 2.1%±4.0%
Table 15: Mean average precision performance for dynamic link forecasting (window-based) over five runs for the discrete-time datasets. Values in parenthesis show the relative change as compared to the average precision performance for dynamic link prediction (batch-based).
Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer $\mu\pm\sigma$
UN V. 52.6( $\downarrow$22.4%) 49.6( $\downarrow$26.8%) 49.7( $\downarrow$24.8%) 52.7( $\uparrow$0.9%) 52.4( $\uparrow$3.4%) 84.2( $\downarrow$0.7%) 52.4( $\downarrow$1.7%) 54.0( $\uparrow$0.1%) 62.4( $\uparrow$4.0%) 9.4%±11.6%
US L. 46.0( $\downarrow$4.5%) 62.5( $\downarrow$14.5%) 58.6( $\downarrow$27.8%) 71.0( $\downarrow$0.3%) 80.7( $\downarrow$0.1%) 63.2( $\downarrow$0.2%) 77.5( $\uparrow$0.1%) 86.5( $\uparrow$0.6%) 86.1( $\uparrow$0.4%) 5.4%±9.6%
UN Tr. 52.7( $\downarrow$10.6%) 49.4( $\downarrow$16.6%) 53.2( $\downarrow$9.7%) 58.9( $\uparrow$1.7%) 59.2( $\uparrow$2.4%) 79.0( $\downarrow$2.6%) 57.5( $\uparrow$2.3%) 65.8( $\uparrow$3.3%) 67.1( $\uparrow$4.4%) 6.0%±5.2%
Can. P. 52.1( $\downarrow$1.3%) 61.0( $\downarrow$1.3%) 69.9( $\uparrow$2.0%) 70.8( $\uparrow$4.5%) 68.3( $\uparrow$6.9%) 59.4( $\downarrow$6.8%) 68.2( $\uparrow$6.2%) 80.9( $\uparrow$4.8%) 83.2( $\downarrow$14.3%) 5.3%±4.0%
Flights 65.2( $\downarrow$2.2%) 63.9( $\downarrow$4.3%) 68.3( $\downarrow$0.0%) 73.5( $\uparrow$1.1%) 65.5( $\uparrow$0.8%) 70.4( $\downarrow$0.2%) 71.0( $\uparrow$0.4%) 71.9( $\uparrow$1.0%) 69.1( $\uparrow$0.7%) 1.2%±1.3%
Cont. 94.0( $\downarrow$0.2%) 95.8( $\uparrow$0.3%) 97.0( $\uparrow$0.8%) 96.8( $\uparrow$0.8%) 88.2( $\uparrow$4.4%) 89.4( $\uparrow$0.6%) 96.6( $\uparrow$1.8%) 95.7( $\uparrow$1.6%) 98.3( $\uparrow$0.6%) 1.2%±1.3%
$\mu\pm\sigma$ 6.9%±8.5% 10.6%±10.4% 10.8%±12.5% 1.6%±1.5% 3.0%±2.5% 1.8%±2.6% 2.1%±2.2% 1.9%±1.8% 4.1%±5.3%
Table 16: Mean average precision performance over five runs for the test set of the continuous-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. Enron 80.8 ± 5.3 78.3 ± 2.3 68.6 ± 5.6 71.7 ± 1.2 76.0 ± 0.7 81.1 ± 0.0 78.2 ± 2.9 89.8 ± 0.4 85.3 ± 0.6
UCI 87.0 ± 1.9 59.5 ± 2.3 69.2 ± 1.0 64.4 ± 1.1 64.0 ± 0.7 68.6 ± 0.0 65.2 ± 0.8 85.3 ± 0.6 80.5 ± 0.9
MOOC 82.4 ± 4.9 76.9 ± 4.3 86.3 ± 2.3 82.7 ± 0.7 72.3 ± 1.3 59.0 ± 0.0 74.9 ± 0.7 74.4 ± 0.6 82.1 ± 8.7
Wiki. 84.1 ± 0.5 80.9 ± 0.4 88.5 ± 0.4 87.5 ± 0.2 75.1 ± 1.0 73.3 ± 0.0 89.2 ± 0.3 90.8 ± 0.2 83.1 ± 1.2
LastFM 76.7 ± 0.6 69.4 ± 1.8 78.8 ± 3.5 72.1 ± 0.8 68.2 ± 0.5 73.4 ± 0.0 70.3 ± 6.5 70.0 ± 1.1 80.5 ± 0.9
Myket 64.5 ± 1.8 63.1 ± 1.5 62.8 ± 2.2 57.9 ± 0.4 46.6 ± 0.2 51.9 ± 0.0 58.9 ± 2.6 60.0 ± 0.2 46.1 ± 1.7
Social 89.4 ± 4.7 91.9 ± 1.0 93.9 ± 1.7 95.0 ± 0.3 85.6 ± 0.1 79.7 ± 0.0 96.1 ± 0.4 95.8 ± 0.2 97.7 ± 0.1
Reddit 80.1 ± 0.3 79.2 ± 0.9 80.6 ± 0.6 78.6 ± 1.0 81.3 ± 0.4 73.5 ± 0.0 76.5 ± 0.6 77.5 ± 0.5 82.8 ± 0.8
Avg. Rank 3.5 5.9 3.8 5.2 7.2 6.9 5.1 4.0 3.4
Pred. Enron 72.1 ± 3.0 69.7 ± 3.7 65.3 ± 3.2 63.2 ± 0.5 66.0 ± 0.5 76.9 ± 0.0 70.2 ± 3.4 82.3 ± 0.6 76.4 ± 0.4
UCI 81.1 ± 3.3 49.0 ± 4.5 71.2 ± 1.1 68.6 ± 1.1 65.1 ± 0.6 65.0 ± 0.0 69.0 ± 0.8 85.9 ± 0.5 80.7 ± 1.1
MOOC 83.4 ± 4.3 77.1 ± 3.8 87.0 ± 2.1 84.5 ± 0.7 73.5 ± 1.0 60.7 ± 0.0 78.5 ± 0.5 77.9 ± 0.8 82.4 ± 9.3
Wiki. 84.1 ± 0.5 80.9 ± 0.3 88.8 ± 0.4 87.9 ± 0.2 75.0 ± 1.1 73.1 ± 0.0 89.5 ± 0.3 91.2 ± 0.2 83.1 ± 1.1
LastFM 77.6 ± 0.6 71.4 ± 1.7 80.3 ± 3.2 75.0 ± 0.7 69.8 ± 0.5 73.2 ± 0.0 71.6 ± 6.1 74.1 ± 1.3 81.1 ± 0.9
Myket 63.4 ± 1.7 62.5 ± 1.4 62.1 ± 2.3 57.1 ± 0.4 45.1 ± 0.2 51.3 ± 0.0 58.3 ± 2.2 59.1 ± 0.2 44.7 ± 1.6
Social 88.3 ± 4.8 91.6 ± 0.7 93.2 ± 2.4 94.8 ± 0.3 86.2 ± 0.2 80.6 ± 0.0 96.2 ± 0.2 95.4 ± 0.1 97.3 ± 0.1
Reddit 80.1 ± 0.3 79.2 ± 0.9 80.5 ± 0.5 78.6 ± 1.0 81.1 ± 0.4 73.7 ± 0.0 76.5 ± 0.6 77.5 ± 0.5 82.7 ± 0.8
Avg. Rank 3.6 6.2 3.6 5.1 7.1 7.4 4.9 3.5 3.5
Table 17: Mean average precision performance over five runs for the test set of the discrete-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. UN V. 52.6 ± 1.8 49.6 ± 1.9 49.7 ± 3.9 52.7 ± 2.6 52.4 ± 2.0 84.2 ± 0.0 52.4 ± 0.9 54.0 ± 1.4 62.4 ± 1.7
US L. 46.0 ± 0.9 62.5 ± 3.6 58.6 ± 2.4 71.0 ± 8.9 80.7 ± 3.7 63.2 ± 0.0 77.5 ± 4.3 86.5 ± 1.9 86.1 ± 1.0
UN Tr. 52.7 ± 3.0 49.4 ± 0.9 53.2 ± 1.5 58.9 ± 2.7 59.2 ± 1.7 79.0 ± 0.0 57.5 ± 1.9 65.8 ± 1.9 67.1 ± 2.7
Can. P. 52.1 ± 0.5 61.0 ± 7.6 69.9 ± 0.8 70.8 ± 1.6 68.3 ± 2.3 59.4 ± 0.0 68.2 ± 1.6 80.9 ± 0.5 83.2 ± 2.9
Flights 65.2 ± 2.7 63.9 ± 2.8 68.3 ± 2.2 73.5 ± 0.3 65.5 ± 1.3 70.4 ± 0.0 71.0 ± 0.4 71.9 ± 0.8 69.1 ± 1.8
Cont. 94.0 ± 2.6 95.8 ± 0.4 97.0 ± 0.5 96.8 ± 0.2 88.2 ± 0.2 89.4 ± 0.0 96.6 ± 0.4 95.7 ± 0.2 98.3 ± 0.0
Avg. Rank 7.7 7.7 5.8 3.5 5.7 4.7 5.0 2.8 2.2
Pred. UN V. 67.8 ± 1.9 67.8 ± 1.7 66.1 ± 3.9 52.3 ± 2.5 50.7 ± 1.4 84.8 ± 0.0 53.3 ± 1.3 53.9 ± 1.7 59.9 ± 1.4
US L. 48.2 ± 1.0 73.1 ± 2.2 81.2 ± 2.1 71.2 ± 8.2 80.8 ± 3.5 63.3 ± 0.0 77.4 ± 4.5 86.0 ± 2.0 85.8 ± 1.0
UN Tr. 58.9 ± 3.1 59.3 ± 1.8 58.9 ± 1.5 57.9 ± 2.4 57.9 ± 2.1 81.1 ± 0.0 56.2 ± 1.5 63.8 ± 1.6 64.3 ± 2.2
Can. P. 52.8 ± 0.5 61.8 ± 1.1 68.5 ± 2.1 67.7 ± 1.7 63.9 ± 1.3 63.8 ± 0.0 64.2 ± 2.0 77.2 ± 0.4 97.1 ± 0.7
Flights 66.7 ± 3.3 66.8 ± 1.6 68.3 ± 1.8 72.7 ± 0.2 65.0 ± 1.3 70.5 ± 0.0 70.8 ± 0.5 71.2 ± 0.7 68.6 ± 1.7
Cont. 94.2 ± 1.3 95.5 ± 0.3 96.3 ± 1.1 96.0 ± 0.3 84.5 ± 0.2 88.8 ± 0.0 95.0 ± 0.6 94.2 ± 0.1 97.7 ± 0.1
Avg. Rank 6.5 5.3 4.0 5.0 7.5 4.8 5.7 3.5 2.7
G.4 AUC-ROC Scores over Time
We present the AUC-ROC scores across the test time for all models and continuous-time datasets with an average relative change in performance (rightmost column in Table 3) larger than one, i.e. Enron (Figure 9), UCI (Figure 10), MOOC (Figure 11), and LastFM (Figure 12). We observe that the patterns seen in Figure 4 for the Enron and MOOC dataset extend to all models. For UCI in Figure 10, we can see the AUC-ROC scores for many time windows and comparably fewer batches. This is because the test set for this dataset contains 15% of all temporal edges but two thirds of the whole observation time.
Figure 12 showing the LastFM dataset contains several outliers in the first half of the test time where the performance is better compared to the average. These outliers correspond to the peaks in activity shown in Figure 5. Similar to what can be observed for the one anomaly in the Enron dataset, these peaks in activity lead to many batches but only a small number of time windows. This is because these bursts occur in a short period of time, leading to an overrepresentation of these outliers using the batch-based approach. At the end of the LastFM dataset, fewer edges occur leading to a less reliable model performance which is underrepresented in the batch-based evaluation due to the fewer number of edges.
Figure 9: Model performance of Enron dataset over time which is visualised by the AUC-ROC scores for each batch and time window individually. The time (x-axis) is limited to the test set.
Figure 10: Model performance of UCI over time.
Figure 11: Model performance of MOOC over time.
Figure 12: Model performance of LastFM over time.
G.5 Investigating Information Leakage
The problem of information leakage can give memory-based models an unfair advantage during evaluation because these models obtain more information than non-memory-based models due to this leakage. However, whether this leaked information is useful for making better predictions depends on the data. We assume that the order of the edges within each snapshot plays an important role in many cases where the leaked information is useful. The data could, for example, retain an implicit temporal ordering if the data was obtained by coarse-graining a continuous-time graph into snapshots. The edges inside the snapshots could also be sorted by some ID, like the node ID of the source node. This would mean that all occurrences of each node as a source node would be consecutive, making it very unlikely that each node appears again as a source node after it was already seen in a previous batch of the same snapshot.
In Figure 13, we show the AUC-ROC scores of each batch individually in the evaluated order and highlight the borders of each snapshot with grey dashed lines. We observe upwards trends in the red regression lines for the US Legislators (DyRep, TGN), the UN Trade (all models), and the Flights (JODIE, DyRep) datasets. This suggests that these datasets contain edges in an order that simplifies the prediction for the indicated models given knowledge from previous batches. Thus, shuffling the data within each snapshot should destroy the useful leaked information and lead to a drop in performance. We report the changes in performance between (i) the models that are evaluated based on edges in the order found in the dataset and (ii) the models evaluated on the datasets with the edges inside each snapshot shuffled in Table 18.
The tables contain the relative change in performance for both evaluation approaches: the batch-based link prediction and our window-based link forecasting. We observe that the performance of the batch-based approach drops for the above-identified dataset and model combinations, suggesting that the edges are indeed ordered in such a way that the earlier batches in a snapshot provide information about later batches, making the prediction easier. Our time-window-based approach prevents this information leakage. Thus, the performance after shuffling is similar to the performance without shuffling because no information, whether useful or not, is leaked.
Table 18: Relative change of the performance between test sets with the order of edges within snapshots as provided by the discrete-time datasets and test sets with edges shuffled within each snapshot. We compare the evaluation approaches of our window-based link forecasting with batch-based link prediction on all memory-based models.
(a) AUC-ROC scores
Model JODIE DyRep TGN
Eval Forec. Pred. Forec. Pred. Forec. Pred.
UN V. $\downarrow$1.6$\downarrow$20.0$\downarrow$7.1$\downarrow$21.4$\downarrow$1.7$\downarrow$15.3%
US L. $\downarrow$0.1$\downarrow$3.5$\downarrow$0.1$\downarrow$10.9$\downarrow$0.6$\downarrow$10.4%
UN Tr. $\downarrow$0.2$\downarrow$13.8$\uparrow$0.1$\downarrow$15.5$\downarrow$3.5$\downarrow$14.2%
Can. P. $\downarrow$0.3$\uparrow$4.3$\downarrow$4.4$\downarrow$0.1$\downarrow$14.0$\downarrow$10.1%
Flights $\uparrow$1.8$\downarrow$1.4$\uparrow$1.3$\downarrow$2.0$\uparrow$0.3$\uparrow$0.1%
Cont. $\downarrow$0.9$\downarrow$1.2$\uparrow$0.1$\downarrow$1.1$\downarrow$0.1$\downarrow$0.2%
$\mu$ 0.8% 7.4% 2.2% 8.5% 3.4% 8.4%
(b) Average precision scores
Model JODIE DyRep TGN
Eval Forec. Pred. Forec. Pred. Forec. Pred.
UN V. $\downarrow$4.2$\downarrow$18.4$\downarrow$6.1$\downarrow$20.5$\downarrow$1.5$\downarrow$13.1%
US L. $\downarrow$0.0$\downarrow$2.2$\downarrow$0.1$\downarrow$11.0$\downarrow$1.5$\downarrow$10.5%
UN Tr. $\downarrow$0.2$\downarrow$10.0$\uparrow$0.9$\downarrow$11.4$\downarrow$3.3$\downarrow$9.9%
Can. P. $\downarrow$0.2$\uparrow$3.1$\downarrow$5.3$\downarrow$0.1$\downarrow$17.6$\downarrow$9.1%
Flights $\uparrow$1.0$\downarrow$2.0$\uparrow$0.9$\downarrow$3.7$\uparrow$0.4$\uparrow$0.7%
Cont. $\downarrow$2.3$\downarrow$3.1$\downarrow$0.2$\downarrow$2.4$\downarrow$0.2$\downarrow$0.6%
$\mu$ 1.3% 6.5% 2.2% 8.2% 4.1% 7.3%
(a) UN Vote
(b) US Legislators
(c) UN Trade
(d) Canadian Parliament
(e) Flights
Figure 13: AUC-ROC score for each batch individually in the order the edges are found in each discrete-time dataset (except Contacts since it has too many snapshots for visualisation). Grey dashed vertical lines separate snapshots, i.e. all points within the same dashed lines correspond to the performances of batches of the same snapshot. The red line represents the best-fit line obtained by performing linear regression on the points within each snapshot.